### v1

Abstract
The welfare of a dynasty that consists of identical Bernoullian utilities subject to Cobb-Douglas production with constant labor is maximized. We blend two discounting schemes to avoid the objective function from exploding to infinity: one is the compound discounting of future utilities and the other is the binary discounting which undiscounts utilities before the planning horizon but fully discounts those that lie beyond. After presenting a general solution to the optimal schedule of consumption, we study the planning solution analytically and also numerically, to discover that the optimal planning horizon is not necessarily an infinity, even under the case of zero compound discounting. Furthermore, we study the inequality of consumptions among generations with respect to the compound discount rates and planning horizons.
[…]
X00 Y00
Keywords:
Dynamic Programming Sustainability Optimal Population Repugnant Conclusion Social Discounting Intergenerational Equity
journal:
1 Introduction
This study is concerned with population axiology from a dynamical perspective. Greaves (2017) delineates population axiology as the study of the conditions under which one state of affairs is better than another, when the states of affairs in question may differ over the numbers and the identities of the persons. There are several approaches to population axiology; one most commonly employed in economics is the total view (utilitarianism) that evaluates populations according to the total amount of well-being that they contain. Among the variety of principles of aggregating potential well-beings, a critical-level utilitarian (e.g., Blackorby et al, 2005) particularly seeks for a state $x$ to maximize the population value $\mathcal{\mathcal{V}}=\sum_{i\in n(x)}\left({\Upsilon}_{i}(x)-\kappa\right)$, where $n(x)$ is the set of individuals, ${\Upsilon}_{i}(x)$ is the lifetime utility of an individual $i$, and $\kappa\geq 0$ is the critical level.
[…]
Under classical utilitarianism ($\kappa=0$), $s=\bar{c}>0$ is called the well-being subsistence (Dasgupta, 2019) or consumption for neutrality, since ${\Upsilon}(s)=0$. For any consumption $c_{i}<s$, utility level of an individual $i$ is negative, meaning that $i$’s life is not worth living. Hence, it must be that $c_{i}>s$ so the lifetime utility of all individuals are positive. Note however that under critical-level utilitarianism, an individual $i$’s life contributes positively to the value of a population if $\bar{c}<c_{i}$, while negatively if $s<c_{i}<\bar{c}$, even the person leads a life worth living. For analytical purposes, we employ ${\Upsilon}(c_{i})=\log c_{i}+\kappa$ with a $\kappa$ that the well-being subsistence $s=e^{-\kappa}$ is sufficiently close to zero. The population value of critical-level utilitarianism will then be $\mathcal{\mathcal{V}}=\sum_{i=1}^{n}\log c_{i}$ where $c_{i}\in[e^{-\kappa},\infty)$ and $\bar{c}=1$, which we hereafter seek to maximize, with respect to $c_{i}$ and $n$.
As we recall our primary question, let us introduce descrete time in population ethics while assuming, for sake of simplicity, that periodical number of people is fixed to one unit; that is, all (representative) individuals are identified by the period $t$ they live in. [^2] We may therefore call $t$ a generation. The population ethics problem then becomes a generation ethics problem. The critical-level utilitarian population value function now becomes $\mathcal{\mathcal{V}}=\sum_{t=0}^{n}\log c_{t}$. We note that this type of value function (based on Bernoullian utility) is widely employed in the previous Ramsey-Cass-Koopmans type of dynastic optimization models (e.g., Nordhaus, 1992; Stern, 2007), except that the future generations are discounted according to the periodical deviation from the present, and typically, the number of potential generations $n$ is left unquestioned (and hence, set to infinity). In contrast, potential generations will be accounted equally whilst the number of generations will be subject to maximization, in this study.
The remainder of the paper proceeds as follows. In the following section, we first specify the model of a production economy based on Cobb-Douglas technology from a dynamical perspective. We then specify our primary question in the form of a finite horizon dynamical programming problem where the objective is to maximize the sum of utilities given the abovementioned model of production, and derive a general solution function. In Section 3 we introduce two broad settings of parameters to simplify the general solution function. The first one is the AK setting where the output elasticity of capital is set to unity. The second one is the ZD (zero discounting) setting where the discount factor is set to unity. We also study the peculiarity of the two models in regard to intergenerational equity and optimal initial action of consumption. Section 4 concludes the paper.
[…]
Figure 1: Contribution to the population value (vertical axis) under critical-level utilitarianism. Left: The critical-level consumption $\bar{c}$ contributes none to the population value, i.e., ${\Upsilon}(\bar{c})-\kappa=0$. Consumption level $c^{*}$ maximizes the critical-level utilitarian population value. Right: Well-being subsistence $s$ bears zero lifetime utility (neutrality), i.e., ${\Upsilon}(s)=0$.
2 The model
2.1 The problem of a dynasity
Let us begin by postulating a two-factor Cobb-Douglas production function of constant returns to scale, as follows:
$Y_{t}={A}(K_{t})^{\theta}(L_{t})^{1-\theta}=\frac{{B}}{\theta^{\theta}(1-\theta)^{1-\theta}}(K_{t})^{\theta}(L_{t})^{1-\theta}$ (1)
where $Y_{t}$ denotes total output of the economy, $K_{t}$ denotes capital, and $L_{t}$ denotes labor, all effective in period $t$. As for the parameters concerned, $\theta<1$ denotes the output elasticity of capital, ${A}>0$ denotes the level of technology termed productivity, and ${B}={A}\Theta$, where $\Theta=\theta^{\theta}(1-\theta)^{1-\theta}$, denotes the cost function-based productivity, which we otherwise call the dual productivity.333 See Appendix 1 for more details.
Below describes the breakdown of total output into investment, depreciation of capital, and final consumption:
$Y_{t}=K_{t+1}-K_{t}+\delta K_{t}+C_{t}$ (2)
Here, $\delta\leq 1$ denotes depreciation factor of capital, and $C_{t}$ denotes consumption in period $t$. Combining equations (1) and (2), and dividing both sides by $L_{t}=L$ which we assume to be constant over time, leads to the following intertemporal dynamics of capital intensity:
$y_{t}={A}(k_{t})^{\theta}=k_{t+1}-(1-\delta)k_{t}+c_{t}$ (3)
where $k_{t}=K_{t}/L$ and $k_{t+1}=K_{t+1}/L$ denote capital intensities in period $t$ and $t+1$, respectively. Also, $y_{t}=Y_{t}/L$ and $c_{t}=C_{t}/L$ denote per caput output and consumption, respectively, in period $t$.
[…]
2.2 Finite horizon dynamic programming
Our primary problem here is to solve for an optimal consumption path given any planning horizon. The Bellman equation of the problem is of the following:555 Note that the objective function of the primary problem (2.1) is given at $t=0$, i.e., $\mathcal{V}[n]=\mathcal{V}_{0}[k_{0};n]$.
$\mathcal{\mathcal{V}}_{t}\left[k_{t};\ n\right]=\max_{c_{t}}\left(\log c_{t}+\beta\mathcal{\mathcal{V}}_{t+1}\left[k_{t+1}={A}(k_{t})^{\theta}-c_{t};\ n\right]\right)$ (4)
The optimal trajectory of the state variable $k_{t}$ is given, by Lemma 2 which we append to Appendix 2 with proofs, as follows:
$k_{t}^{*}[n]=(k_{0})^{\theta^{t}}\prod_{i=1}^{t}\left(\frac{S_{{n}-i}}{S_{{n}-i+1}}{A}\beta\theta\right)^{\theta^{t-i}}=(k_{0})^{\theta^{t}}\prod_{i=1}^{t}\left(\frac{S_{{n}-t+i-1}}{S_{{n}-t+i}}{A}\beta\theta\right)^{\theta^{i-1}}$ (5)
where $S_{\ell}$ is defined as follows:
$S_{\ell}\equiv\sum_{i=0}^{\ell}(\beta\theta)^{i}=1+(\beta\theta)+(\beta\theta)^{2}+\cdots+(\beta\theta)^{\ell}=\frac{1-(\beta\theta)^{\ell+1}}{1-\beta\theta}$ (6)
The following optimal trajectory of consumption is obtained by (5), and (21) which must be true by the proof of Lemma 1.
$c_{t}^{*}[n]=\frac{{A}\left(k_{t}^{*}[n]\right)^{\theta}}{S_{n-1}}=\frac{{A}(k_{0})^{\theta^{t+1}}}{S_{{n}-t}}\prod_{i=1}^{t}\left(\frac{S_{{n}-t+i-1}}{S_{{n}-t+i}}{A}\beta\theta\right)^{\theta^{i}}$ (7)
With formula (7) the population value function is given as follows:
$\mathcal{V}[n]$ $=\sum_{t=0}^{n}\beta^{t}\log c_{t}^{*}[n]$
$=\log\left(\frac{{A}(k_{0})^{\theta}}{S_{n}}\right)+\sum_{t=1}^{n}\beta^{t}\log\left(\frac{{A}(k_{0})^{\theta^{t+1}}}{S_{{n}-t}}\prod_{i=1}^{t}\left(\frac{S_{{n}-t+i-1}}{S_{{n}-t+i}}{A}\beta\theta\right)^{\theta^{i}}\right)$ (8)
We hereafter seek to maximize this function with respect to the planning horizon $n$.
Our approach to the analysis of population value (8) will be numerical rather than analytical, because the derivative for $\mathcal{V}[n]$ with respect to $n$ seems uninformative. As it turns out in the following section, the optimal tragectory of contributions $\log c_{t}^{*}[n]$ and the population value $\mathcal{V}[n]$, for any given planning horizon $n$, becomes manageable under $\theta=1$ (known as the AK setting) and $\beta<1$. We also find that the trajectory of contributions $\log c_{t}^{*}[\infty]$ and the population value $\mathcal{V}[\infty]$ for an infinite planning horizon, is evaluable under $\beta\theta<1$. We therefore base our study of zero discounting ($\beta=1$) on this setting (i.e., $\beta\theta<1$, which indicates $\theta<1$, a Cobb-Douglas model). In the following section, we delve into the abovementioned two broad settings of parameters, namely, AK production with future discounting ($\theta=1$ and $\beta<1$) which we term as AK settings, and Cobb-Douglas production without future discounting ($\theta<1$ and $\beta=1$) which we term as ZD settings.
Table 1 summarizes the parameter settings chosen for numerical examinations. Note that cases I–IV correspond to AK settings whose solution paths are characterized by the sign of $\log({A}\beta)$, whereas cases V–VII correspond to ZD settings whose solution paths are characterized by the sign of $\log(A\Theta)$. The final case VIII corresponds to the parameter settings that exemplifies $\log({A}\beta)=\log(A\Theta)=0$.
Table 1: Parameters applied in various cases.
case ${A}$ $\beta$ $\theta$ $\log({A}\beta)$ $\log({A\Theta})$ $n^{*}$ $\mathcal{V}[n^{*}]$ $\mathcal{V}[\infty]$
I $1.012$ $0.992$ $1$ $+$ $\infty$ $84.675$
II $1.01$ $0.992$ $1$ $+$ 95 $59.965$
III *1 $0.992$ $1$ 00 73 $55.627$
IV $1.005$ $0.992$ $1$- 58 $50.979$
V $1.05$ $1$ $0.992$ $+$ $\infty$ +$\infty$+ ∞
VI $1.2$ $1$2 00 (281) $41.688$
VII $1.05$ $1$ $0.991$- 117 $\infty$- ∞
VIII $1$ $1$ $1$ 00 00 54 $55.182$$\infty$- ∞
  • Note:
    For all cases I–VIII, $k_{0}=150$.
  • *1
    ${A}=1/\beta\approx 1.008065$.
  • *2
    $\theta\approx 0.955392$ where 1.2($1{1.2}(1-\theta)^{1-\theta}\theta^{\theta}=$ 1.
3 Analysis
3.1 AK setting
AK model is one of the simplest yet fundamental models of endogenous growth, which was formally developed by Frankel (1962). Here, we take this production model to study the population value function that discounts future generations to see whether if $n\to\infty$ is an optimal policy. The optimal consumption path (7) for AK setting with future discounting ($\theta=1$, $\beta<1$) may be specified as follows:
$c_{t}^{*}[n]=\left(\frac{{A}k_{0}}{S_{{n}-t}}\right)\left(\frac{S_{{n}-1}}{S_{n}}{A}\beta\right)\left(\frac{S_{{n}-2}}{S_{{n}-1}}{A}\beta\right)\cdots\left(\frac{S_{{n}-t}}{S_{{n}-t+1}}{A}\beta\right)=\frac{({A}\beta)^{t}{A}k_{0}}{S_{n}}$ (9)
The population value function (8), therefore, becomes:
$\mathcal{V}[n]$ $=\sum_{t=0}^{n}\beta^{t}\log\left(\frac{({A}\beta)^{t}{A}k_{0}}{S_{n}=1+\beta+\cdots+\beta^{n}}\right)$
$=\frac{\beta-((1-\beta){n}+1)\beta^{{n}+1}}{(1-\beta)^{2}}\log({A}\beta)+\frac{1-\beta^{{n}+1}}{1-\beta}\log\left(\frac{1-\beta}{1-\beta^{{n}+1}}{A}k_{0}\right)$ (10)
For sake of the analysis let us take the derivative with respect to ${n}$.
$\frac{\mathrm{d}\mathcal{V}[n]}{\mathrm{d}n}=\lambda\left(\gamma-{n}\log({A}\beta)+\log\left(1-\beta^{n+1}\right)\right)$ (11)
where $\lambda$ and $\gamma$, that are both independent of $n$, are given as follows:
$\lambda=\frac{\beta^{n}\log\beta^{\beta}}{1-\beta}>0,$ $\gamma=1-\frac{\log({A}\beta)^{1-\beta+\log\beta}}{\log\beta^{1-\beta}}-\log({A}(1-\beta)k_{0})$
In order to study the derivative sign of (11), we consider the following two functions whose equality convey the first order conditions, i.e., $f[n]=g[n]$.
$f[n]=-\gamma+n\log({A}\beta),$ $g[n]=\log\left(1-\beta^{n+1}\right)$ (12)
Before we proceed, let us examine the possible range of the parameter $A$. Described below is the evaluation of the marginal product of capital (MPK) of our production function (1) net of depreciation, which should coincide with the real interest rate $\rho>0$, viz.,
$\frac{\partial Y_{t}}{\partial K_{t}}-\delta={A}\theta\left(\frac{K_{t}}{L_{t}}\right)^{\theta-1}-\delta=\rho>0$ (13)
Then, we know that AK models with complete capital depreciation $\theta=\delta=1$ amounts to ${A}=1+\rho$, and therefore ${A}>1$ is a relevant assumption in this setting.666 Note also that $A\beta=(\delta+\rho)\beta$ indicates the discrepancy between the gross rate of interest and the rate of generational preference of the population.
Figure 2 (left) depicts functions $f[n]$ and $g[n]$ under different parameters for cases I–IV. Clearly, $f[n]$ is a linear function whose slope is $\log({A}\beta)$, and $g[n]$ is monotonously increasing and approaching zero, i.e., $g[\infty]=0$. Here, we fix the discount factor $\beta$, and differentiate the productivity ${A}$ in four different levels. As long as ${A}\leq 1/\beta$, so that $\log({A}\beta)\leq 0$, i.e., the slope is zero or negative, $f[n]$ will intersect with $g[n]$ at a single point and the population value function $\mathcal{V}[n]$ has a single peak. If ${A}>1/\beta$ so that $f[n]$ has a positive slope, then $f[n]$ and $g[n]$ could either intersect with two points where the population value $\mathcal{V}[n]$ would rise, fall and then rise again, or never intersect so that the population value would be rising indefinitely, with respect to the planning horizon $n$. Figure 2 (right) depicts the population value functions for cases I–IV.
Figure 2: Left: Solid plots show $g[n]$ defined in (12) for cases I–IV with obvious correspondences, i.e., the steepest one corresponds to case I, whereas case IV corresponds to the one with a negative slope. Dashed plot shows $f[n]$ also defined in (12). Right: The plots show population value paths $\mathcal{V}[n]$ for cases I–IV with obvious correspondences.
To visualize the optimal trajectory of the variables in various situations, we specify them here in the form of functionals. By virtue of (9), the optimal trajectory of undiscounted contribution to the population value becomes linear with respect to $t$, as follows:
$\log c_{t}^{*}[n]=\log\left(\frac{1-\beta}{1-\beta^{n+1}}{A}k_{0}\right)+t\log({A}\beta)$ (14)
By virtue of (21), the optimal trajectory of capital intensity becomes as follows:
$k_{t}^{*}[n]=\frac{S_{n-t}}{{A}}c_{t}^{*}[n]=\frac{1-\beta^{n+1-t}}{1-\beta^{n+1}}({A}\beta)^{t}k_{0}$ (15)
Figure 3 displays the population value function $\mathcal{V}[n]$, specified as (10), for cases I, III, and IV, on the top. [^7] These figures are reiterations from Figure 2 (right). Note that, as long as $\beta<1$, the population value function will always converge to some finite value, regardless of the parameter settings besides $\beta$. As we let $n\to\infty$ in (10), we have what follows:
$\mathcal{V}[\infty]=\frac{\log\left({A}(1-\beta)^{1-\beta}\beta^{\beta}\right)}{(1-\beta)^{2}}+\frac{\log(k_{0})}{1-\beta}$ (16)
By comparing the population values at the numerical solution of the first order condition $f[n]=g[n]$ in reference to (12), with $\mathcal{V}[\infty]$ in reference to (16), we know the optimal planning horizon $n^{*}$, which we display in Table 1 for cases I–IV. At the middle row of Figure 3, optimal trajectories of undiscounted contribution to the population value, according to (14), with respect to planning horizons $n=200,400,600$, are displayed for sample I, III, and IV from left to right. Similarly at the bottom row of Figure 3, optimal trajectories of capital intensity, according to (15), with respect to planning horizons $n=200,400,600$, are displayed for cases I, III, and IV from left to right.
Figure 3: Population value function (top), trajectories of undiscounted contribution (middle) and capital intensity (bottom), for AK model cases I (left), III (center), and IV (right).
3.2 ZD setting
Here, we study zero discounting $\beta=1$ under Cobb-Douglas production $\theta<1$. The condition that $\beta\theta<1$ nonetheless allows us to evaluate the key summation as follows:
$S_{\infty-\tau}=\lim_{n\to\infty}\frac{1-(\beta\theta)^{n-\tau+1}}{1-\beta\theta}=\frac{1}{1-\beta\theta}$
As we apply the above and $n\to\infty$ into (7), we have the following:
$\log c_{t}^{*}[\infty]$ $=\log\left(\frac{{A}}{S_{{\infty}-t}}\right)+\theta\left(\sum_{i=1}^{t}\theta^{t-i}\log\left(\frac{S_{{\infty}-i}}{S_{{\infty}-i+1}}{A}\beta\theta\right)+\theta^{t}\log k_{0}\right)$
$=\log\left({{A}}{(1-\beta\theta)}\right)+\left(\theta^{t-1}+\cdots+\theta+1\right)\log({A}\beta\theta)^{\theta}+\theta^{t+1}\log k_{0}$
$=\frac{\log\left({A}(1-\beta\theta)^{1-\theta}(\beta\theta)^{\theta}\right)}{1-\theta}+\theta^{t}\left(\log(k_{0})^{\theta}-\log({A}\beta\theta)^{\frac{\theta}{1-\theta}}\right)$ (17)
where the optimal trajectory for undiscounted contributions turns out to be geometrically convergent. We can then apply $\beta=1$ to arrive at the following zero discount verstion of the optimal trajectory for undiscounted contributions:
$\log c_{t}^{*}[\infty]=\frac{\log\left({A}(1-\theta)^{1-\theta}\theta^{\theta}\right)}{1-\theta}+\theta^{t}\left(\log(k_{0})^{\theta}-\log({A}\theta)^{\frac{\theta}{1-\theta}}\right)$
The corresponding population value can hence be evaluated by the infinite sum of the undiscounted contributions, i.e.,
$\mathcal{V}[\infty]=\sum_{t=0}^{\infty}\log c_{t}^{*}[\infty]=$ $\frac{\infty\log\left({A}(1-\theta)^{1-\theta}\theta^{\theta}\right)+\log(k_{0})^{\theta}-\log({A}\theta)^{\frac{\theta}{1-\theta}}}{1-\theta}$
$=$ $\begin{cases}+\infty&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}={A}\Theta>1\\ {\frac{\theta}{1-\theta}}\log\left(\frac{1-\theta}{\theta}k_{0}\right)&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}={A}\Theta=1\\ -\infty&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}={A}\Theta<1\end{cases}$
The above result indicates that if ${A}\Theta>1$, then the population value is ever increasing and $n\to\infty$ must be the optimal solution; whereas, if ${A}\Theta<1$, then $n\to\infty$ must not be optimal and the population value must be maximized at a finite horizon $n^{*}\ll\infty$. The case where ${A}\Theta=1$ turns out to be the one that the planning horizon $n$ does not matter (beyond certain length) in maximizing the population value. Case V reflects ${A}\Theta>1$ where $n^{*}=\infty$, case VII reflects ${A}\Theta<1$ where $n^{*}\ll\infty$, and case VI reflects ${A}\Theta=1$ where $n^{*}$ is any number greater than $281$.888 The number is subject to decimals rounding. Figure 4 dipicts the population value function $\mathcal{V}[n]$, the optimum trajectory of undiscounted contribution $\log c_{t}^{*}[n]$ and that of capital intensity $k_{t}^{*}[n]$, given the planning horizon $n$, for sample ZD settings V, VI, and VII.
The final case VIII implies ZD and AK setting ($\theta=\beta=1$), where $A\Theta=A=1+\rho>1$, under complete depreciation, by virtue of (13). Thus, if $\rho>0$ then $A>1$ in which case the population value would explode as described above, and we are left with $n^{*}\to\infty$. If in turn, $A=1+\rho=1$ as the case (VIII) goes, the infinite horizon population value is evaluated as follows:
$\mathcal{V}[\infty]=\lim_{\theta\to 1}{\frac{\theta}{1-\theta}}\log\left(\frac{1-\theta}{\theta}k_{0}\right)=-\infty$
This indicates that there exists a finite optimum horizon $n^{*}\ll\infty$. Note that the population value function in this setting can be specified by zero discounting $(\beta\to 1)$ the AK model’s population value function (10), as follows:
$\mathcal{V}[n]$ $=\lim_{\beta\to 1}\left(\frac{\beta-((1-\beta){n}+1)\beta^{{n}+1}}{(1-\beta)^{2}}\log({A}\beta)+\frac{1-\beta^{{n}+1}}{1-\beta}\log\left(\frac{1-\beta}{1-\beta^{{n}+1}}{A}k_{0}\right)\right)$
$=\lim_{\beta\to 1}\frac{1-\beta^{{n}+1}}{1-\beta}\log\left(\frac{1-\beta}{1-\beta^{{n}+1}}k_{0}\right)=\lim_{\beta\to 1}(n+1)\beta^{n}\log\left(\frac{k_{0}}{(n+1)\beta^{n}}\right)$
$=(n+1)\log\left(\frac{k_{0}}{n+1}\right)$
The proof for the second identity, given $A=1$, is appended to Appendix 3. The third identity is subject to L’Hôpital’s rule. By the first order condition $\frac{\partial\mathcal{V}[n]}{\partial n}=0$, the optimal planning horizon is evaluated as $n^{*}=\exp(-1+\log k_{0})-1\approx 54.182$.
Figure 4: Population value function (top), trajectories of undiscounted contribution (middle) and capital intensity (bottom), for ZD setting cases V (left), VI (center), and VII (right). Note that the population value function of the top-left panel is increasing indefinitely.
3.3 Intergenerational inequality
It seems obvious from Figure 3 (middle row) that consumption inequality among generations is worsened as we expand the planning horizon in the presence of future discounting (except for the knife-edge case III with $\log(A\beta)=0$), while it seems to improve in Figure 4 (middle row), for cases with zero discounting. To check the validity of the conjecture, we prepare Figure 5 to illustrate how Lorenz curve moves in regard to the planning horizon (i.e., $n=200,400,600$) for AK setting cases I, III, IV, and for ZD setting cases V, VI, and VII. The Lorenz curves (in Figure 5) are based on the stream of optimal consumption for a given planning horizon, namely, $c_{t}^{*}[n]$. The level of inequality increases in AK settings as the planning horizon is stretched longer. On the other hand, the level of inequality is relatively insensitive to the planning horizon for zero discounting cases, except for case VI with $\log(A\Theta)=0$ where inequality decreases along with the length of the planning horizon.
Figure 5: Lorenz curves for $n=200$ (solid line), $n=400$ (dotted line), and $n=600$ (dashed line). Top row panels (from left to right) correspond to AK settings I, III, and IV. Second row panels (from left to right) correspond to ZD settings V, VI, and VII.
For further analysis, let us consider Gini Index, a popular measure of inequality, defined on the optimal stream of consumptions $c_{t}^{*}[n]$, as follows:
$\mathcal{G}[n]=\frac{\sum_{t^{\prime}=0}^{n}\sum_{t=0}^{n}\left|c_{t}^{*}[n]-c_{t^{\prime}}^{*}[n]\right|}{2n\sum_{t=0}^{n}c_{t}^{*}[n]}$
Figure 6 illustrates Gini Index $\mathcal{G}$ as a function of planning horizon $n$ in various parameter settings. For both figures the the underlying parameters are fixed at $k_{0}=150$ and $A=1$. The left figure pertains to AK settings ($\theta=1$), while we raise the discount rate from $\beta=0.9$ (solid line) to $\beta=0.99$ (dashed line). As we can see from the (left) panel, AK settings have the tendency to unequalize consumptions between generations as the planning horizon expands, while inequality is mitigated by less discountig the future. In contrast, the right figure indicates that ZD settings have the tendency to equalize consumptions between generations as the planning horizon expands, while inequality is enhanced by the higher output elasticity of capital.
Figure 6: The panels are planning horizon vs Gini Index i.e., $(n,\mathcal{G}[n])$ for AK model cases with different discount rates (left), and zero discounting cases with different output elasticities of capital (right). The parameters setting of the left panel is $\theta=1$, $k_{0}=150$, $A=1$, $\beta=0.9$ (solid line), and $\beta=0.99$ (dashed line). The parameters setting of the right panel is $\beta=1$, $k_{0}=150$, $A=1$, $\theta=0.9$ (solid line), and $\theta=0.95$ (dashed line).
3.4 The initial action
Let us now focus on the initial action of the optimal consumption schedule which we specify, by virtue of (7) or (21), as follows:
$c_{0}^{*}[n]=\frac{{A}(k_{0})^{\theta}}{S_{n}}=\frac{1-\beta\theta}{1-(\beta\theta)^{n}}{A}(k_{0})^{\theta}$ (18)
We take the derivative of (18) and obtain the following result:
$\frac{\mathrm{d}c_{0}^{*}[n]}{\mathrm{d}n}=\frac{A(k_{0})^{\theta}\left(1-\beta\theta\right)(\beta\theta)^{n}\log(\beta\theta)}{\left(1-(\beta\theta)^{n}\right)^{2}}<0$
since, $\beta\theta<1$. In other words, the further we take future generations into account, the more current generation must sacrifice. In any case, the initial action is monotonously decreasing in the planning horizon, and ultimately converge to the following:
$c_{0}^{*}[\infty]=(1-\beta\theta){A}(k_{0})^{\theta}$ (19)
By taking it’s derivative with respect to $\beta$ we obtain:
$\frac{\partial c_{0}^{*}[\infty]}{\partial\beta}=-\theta A(k_{0})^{\theta}<0$
That is to say, the more we respect future generations (by raising $\beta$), the more current generation must sacrefice. From another perspective, we can solve the equation $c_{0}^{*}[n]=s$ for $n$ in regard to (18), where $s$ is the consumption level of subsistence, as follows:
$n=\frac{\log\left((1-\beta\theta)A(k_{0})^{\theta}\right)-\log s}{\log(\beta\theta)}$
This $n$ is the subsistence-proof number of potential generations.
Finally, for AK models, $c_{0}^{*}[\infty]=(1-\beta)Ak_{0}$ by (19), but this value will approach zero if future generations are respected as equally as possible to the current one ($\beta\to 1$). That is to say, the ultimate respect for the future generations in AK models with an infinite planning horizon induces the current generation to sacrefice ultimately, i.e., $c_{0}^{*}[\infty]\to{s}$. Alternatively, if we relax the infinite horizon assumption, initial action for AK models with ultimate respect for the future generations be:
$c_{0}^{*}[n]=\lim_{\beta\to 1}\frac{1-\beta}{1-\beta^{n}}{A}k_{0}=\frac{Ak_{0}}{n}$
and we are left with the AK and ZD version of subsistence-proof number of potential generations $n=Ak_{0}/s$.
4 Concluding Remarks
The extent that we take into account of the entire potential generation upon creating the objective function of the population (i.e., the dynasty) is usually determined by the compounding discounting factor. Notice that the potential number of intrinsic years that are accounted by compound discounting at a rate of $r$ is as much as $\int_{0}^{\infty}e^{-{r}x}\mathrm{d}x=1/r$. As regards to our problem, compound discounting at a rate of $r$ is equivalent to binary discounting at a planning horizon $n=1/r$. That is, for example, a generational discount rate of $1$% amounts to 100 generations. Binary discounting is flexible in a sense that the planning horizon is variable, so that infinity is not precluded from an option. Moreover, it is nondiscriminatory, ensuring that living generations are treated equally. With binary discounting, however, an economic growth model requires a finite horizon dynamic optimization approach.
One important feature of binary discounting is that it hinders future generations that earn below-critical-level utilities. If the phase boundary $\log(A\beta)$ for AK setting is negative, the infinite horizon solution trajectory of utility (less the critical level), is forever declining (as shown in the middle-right panel of Figure 3) and eventually surpasses the level of subsistence. Binary discounting avoids such repugnancy by limiting the longevity of the dynasty. On the other hand, if the phase boundary $\log({A}\Theta)$ for zero (compound) discounting is negative, a finite horizon must be optimal since an infinite horizon objective function approaches a negative infinity. Note also that the bottom utilities that lie at the center of the planning horizon (as shown in the middle-right panel of Figure 4), yield increasingly negative contributions to the population value.
As regards the distribution of consumption among generations, the rate of compound discounting plays the key role. That is, the less the rate of discounting is, the more equal the consumptions become. Moreover, as the planning horizon expands, the distribution of consumption becomes more uneven in the AK setting cases, whereas vice-versa in a zero discounting setting. Our another concern was the optimal initial consumption in light of the well-being of future generations. In any case the initial consumption is decreasing in the planning horizon, and ultimately, it will depend on the value of $\beta\theta$. The optimal initial consumption for infinite planning horizon will decrease as discount rate approaches zero, and under the AK setting, initial consumption itself approaches zero until it reaches the level of subsistence.
Appendix 1
Below we put down a production function and its dual unit cost function:
${Y}=A{K}^{\theta}{L}^{1-\theta},$ $p={B}^{-1}{r}^{\theta}w^{1-\theta}$
where $p$, $r$, and $w$ denote prices for $Y$, $K$, and $L$, respectively. The remaining parameter $A$ is called the productivity and $B$ is called the dual cost function-based productivity. Applying Shephard’s lemma to the dual function leads to the followings.
$\frac{\partial p}{\partial r}=\frac{\theta}{B}\left(\frac{r}{w}\right)^{\theta-1}=\frac{K}{Y},$ $\frac{\partial p}{\partial w}=\frac{1-\theta}{B}\left(\frac{r}{w}\right)^{\theta}=\frac{L}{Y}$
By virtue of these euqations, the marginal product of capital (MPK) can be readily evaluated as follows:
$\text{MPK}=\frac{\partial Y}{\partial K}=A\theta\left(\frac{K}{L}\right)^{\theta-1}$ $=A\theta^{\theta}(1-\theta)^{1-\theta}\left(\frac{r}{w}\right)^{\theta-1}$
$A\theta\left(\frac{K}{L}\right)^{\theta-1}$ $=\frac{Y}{K}\theta=B\left(\frac{r}{w}\right)^{\theta-1}$
where, $r/w\in(0,\infty)$ equals the marginal rate of substitution between capital and labor (MRS). By comparison, we are left with $B=A\theta^{\theta}(1-\theta)^{1-\theta}\equiv A\Theta$. Moreover,
$A=\left(\frac{\text{MPK}}{\Theta}\right)\left(\frac{r}{w}\right)^{\theta-1}$ $B=\text{MPK}\left(\frac{r}{w}\right)^{\theta-1}$
On the other hand, recall the breakdown equation of total output (2) and take the partial derivative as follows:
$\frac{\partial Y_{t}}{\partial K_{t}}=\left(\frac{\partial K_{t+1}}{\partial K_{t}}-1\right)+\delta=\rho+\delta=\text{MPK}$ (20)
where $\rho>0$ denotes the rate of interest. Hence, if $\theta=1$ (AK setting), $A=B=\text{MPK}=\rho+\delta$, and if $\delta=1$ (complete depreciation), it must be the case that $A=B=1+\rho>1$.
According to the study in section 3.1, we note that condition $A\beta<1$ implies the non-increasing of undiscounted contributions (or utilities less the critical level) with respect to generations. The knife-edge case where $A\beta=1$ implies that the undiscounted contributions are constant over generations. Figure 7 (left) dipicts the set of possible values for $(\theta,A)$ that satisfies $A\beta\leq 1$. Also, according to the study in section 3.2, we note that condition $A\Theta<1$ implies the existence of a finite optimal planning horizon. Otherwise, if $A\Theta>1$, the optimal planning horizon will be inifinite. The knife-edge case where $A\Theta=1$ implies that the optimal planning horizon is indeterminate. Figure 7 (right) dipicts the set of possible values for $(\theta,A)$ that satisfies $A\Theta\leq 1$.
[…]
Figure 7: The shaded area (mariginal lines not included) corresponds to the set of parameters $(\theta,A)$ where $B=A\theta^{\theta}(1-\theta)^{1-\theta}<1$
Appendix 2
Lemma 1.
The value function of Bellman equation (4) is of the following:
$\mathcal{V}_{t}\left[k_{t};n\right]$ $=S_{{n}-t}\theta\log(k_{t})+{R}_{t}$
where $S_{n-t}$ is specified by (6) and ${R}_{t}$ is a term that does not depend on $k_{t}$.
Proof.
We show this by induction. First, evaluate the value function at $t={n}$ by (4):
$\mathcal{V}_{n}\left[k_{n};{n}\right]=\max_{c_{n}}\left(\log c_{n}+\beta\mathcal{V}_{n+1}[k_{{n}+1}={A}(k_{n})^{\theta}-c_{n};{n}]\right)$
Since $k_{{n}+1}=0$ by (2.1), $c_{n}={A}(k_{n})^{\theta}$ must be true. Also, $\mathcal{V}_{{n}+1}=0$ must be true for efficiency. Thus, we know that the final maximization is bounded, i.e.,
$\mathcal{V}_{n}\left[k_{n};{n}\right]=\log c_{n}=\log{A}+\theta\log k_{n}$
Because $S_{n-n}=S_{0}=1$ and since ${A}$ is a constant, the proposition holds true for $t={n}$.
Suppose the proposition holds true for $t+1$. Then,
$\mathcal{V}_{t}\left[k_{t};n\right]=$ $\max_{c_{t}}\left(\log c_{t}+\beta\mathcal{V}_{t+1}\left[k_{t+1}={A}(k_{t})^{\theta}-c_{t};n\right]\right)$
$=$ $\max_{c_{t}}\left(\log c_{t}+\beta\left(S_{{n}-t-1}\theta\log\left({A}(k_{t})^{\theta}-c_{t}\right)+{R}_{t+1}\right)\right)$
Below are the corresponding first order condition and its solution:
$\frac{1}{c_{t}}-\frac{(\beta\theta)S_{{n}-t-1}}{{A}(k_{t})^{\theta}-c_{t}}=0,$ or, $c_{t}=\frac{{A}(k_{t})^{\theta}}{S_{{n}-t}}$ (21)
Here, we used $(\beta\theta)S_{{n}-t-1}=\beta\theta+(\beta\theta)^{2}+\cdots+(\beta\theta)^{n-t}=S_{n-t}-1$. By plugging the above solution back into the maximand we arrive at the following result:
$\mathcal{V}_{t}[k_{t};n]$ $=\log\left(\frac{{A}(k_{t})^{\theta}}{S_{{n}-t}}\right)+\beta\left(R_{t+1}+S_{n-t-1}\theta\log\left({A}(k_{t})^{\theta}-\frac{{A}(k_{t})^{\theta}}{S_{{n}-t}}\right)\right)$
$=S_{{n}-t}\theta\log(k_{t})+\left(\beta{R}_{t+1}+S_{{n}-1}\log\left(\frac{{A}}{S_{{n}-t}}\right)+\log(S_{n-1}-1)^{S_{n-1}-1}\right)$
$=S_{{n}-t}\theta\log(k_{t})+R_{t}$
Hence, the lemma. ∎
Lemma 2.
The optimal trajectory of state $k_{t}^{*}[n]$ for Bellman equation (4) is:
$\ln k_{t}^{*}[n]=\sum_{i=1}^{t}\theta^{t-i}\log\left(\frac{S_{{n}-i}}{S_{{n}-i+1}}{A}\beta\theta\right)+\theta^{t}\log k_{0}$
Proof.
We show this by induction. By plugging (21) into (2.1), we obtain:
$k_{t+1}={A}(k_{t})^{\theta}-c_{t}=\frac{S_{{n}-t}-1}{S_{{n}-t}}{A}(k_{t})^{\theta}=\frac{S_{{n}-t-1}}{S_{{n}-t}}{A}\beta\theta(k_{t})^{\theta}$ (22)
As we apply $t=0$ on the above (22) and by taking the logarithm,
$\log k_{1}=\ln\left(\frac{S_{{n}-1}}{S_{n}}{A}\beta\theta\right)+\theta\log k_{0}$
we know that the proposition is true for $t=1$.
Suppose that the proposition is true for $t$. Then, we know by (22) that:
$\log k_{t+1}^{*}[n]$ $=\log\left(\frac{S_{{n}-t-1}}{S_{{n}-t}}{A}\beta\theta\right)+\theta\log k_{t}^{*}[n]$
$=\log\left(\frac{S_{{n}-t-1}}{S_{{n}-t}}{A}\beta\theta\right)+\theta\left(\sum_{i=1}^{t}\theta^{t-i}\log\left(\frac{S_{{n}-i}}{S_{{n}-i+1}}{A}\beta\theta\right)+\theta^{t}\log k_{0}\right)$
$=\sum_{i=1}^{t+1}\theta^{t+1-i}\log\left(\frac{S_{{n}-i}}{S_{{n}-i+1}}{A}\beta\theta\right)+\theta^{t+1}\log k_{0}$
That is, the proposition is true for $t+1$. ∎
Proposition 1.
The optimal consumption trajectory $c_{t}^{*}[n]$ for Bellman equation (4) is:
$\log c_{t}^{*}[n]=\log\left(\frac{{A}}{S_{{n}-t}}\right)+\theta\left(\sum_{i=1}^{t}\theta^{t-i}\log\left(\frac{S_{{n}-i}}{S_{{n}-i+1}}{A}\beta\theta\right)+\theta^{t}\log k_{0}\right)$
Proof.
Obvious from Proposition 2 and (21). ∎
Appendix 3
Let us evaluate the following:
$\lim_{\beta\to 1}\frac{\beta-((1-\beta){n}+1)\beta^{n+1}}{(1-\beta)^{2}}$ $=\lim_{\beta\to 1}\frac{1-(n+1)((1-\beta)n+1)\beta^{n}+n\beta^{n+1}}{-2(1-\beta)}$
$=\lim_{\beta\to 1}\frac{n(n+1)\left(2\beta^{n}-((1-\beta)n+1)\beta^{n-1}\right)}{2}$
$={n(n+1)}/{2}$
where we used the L’Hôpital’s rule twice. Then, we know that
$\lim_{\beta\to 1}\frac{\beta-((1-\beta){n}+1)\beta^{n+1}}{(1-\beta)^{2}}\log(A\beta)=\frac{n(n+1)}{2}\log(A)=0$

### v3

@@
 Abstract
-The welfare of a dynasty that consists of identical Bernoullian utilities subject to Cobb-Douglas production with constant labor is maximized.
-We blend two discounting schemes to avoid the objective function from exploding to infinity: one is the compound discounting of future utilities and the other is the binary discounting which undiscounts utilities before the planning horizon but fully discounts those that lie beyond.
-After presenting a general solution to the optimal schedule of consumption, we study the planning solution analytically and also numerically, to discover that the optimal planning horizon is not necessarily an infinity, even under the case of zero compound discounting.
-Furthermore, we study the inequality of consumptions among generations with respect to the compound discount rates and planning horizons.
+This paper develops a dynamic optimization framework for evaluating the optimal longevity of a dynasty under a critical-level utilitarian social welfare criterion.
+Unlike standard models that assume an infinite planning horizon and apply discounting to future generations, we endogenize the planning horizon itself and examine the welfare-maximizing number of generations in a roundabout production economy with capital accumulation.
+The model incorporates a well-being threshold that must be met for generational existence to be considered ethically justified, providing an alternative to traditional utilitarian aggregation.
+Using finite-horizon dynamic programming, we derive closed-form consumption trajectories and analyze how technological parameters and intertemporal preferences affect the optimal generational span.
+Our results show that the optimal horizon need not be infinite—even under zero discounting—and that intergenerational equity may require limiting dynasty length.
 […]
-X00 Y00
+O21 I30 C61
 Keywords:
-Dynamic Programming Sustainability Optimal Population Repugnant Conclusion Social Discounting Intergenerational Equity
+Critical-level Utilitarianism Dynamic Programming Optimal Population Intergenerational Equity
 journal:
 1 Introduction
-This study is concerned with population axiology from a dynamical perspective.
-Greaves (2017) delineates population axiology as the study of the conditions under which one state of affairs is better than another, when the states of affairs in question may differ over the numbers and the identities of the persons.
-There are several approaches to population axiology; one most commonly employed in economics is the total view (utilitarianism) that evaluates populations according to the total amount of well-being that they contain.
-Among the variety of principles of aggregating potential well-beings, a critical-level utilitarian (e.g., Blackorby et al, 2005) particularly seeks for a state $x$ to maximize the population value $\mathcal{\mathcal{V}}=\sum_{i\in n(x)}\left({\Upsilon}_{i}(x)-\kappa\right)$, where $n(x)$ is the set of individuals, ${\Upsilon}_{i}(x)$ is the lifetime utility of an individual $i$, and $\kappa\geq 0$ is the critical level.
-[…]
-Under classical utilitarianism ($\kappa=0$), $s=\bar{c}>0$ is called the well-being subsistence (Dasgupta, 2019) or consumption for neutrality, since ${\Upsilon}(s)=0$.
-For any consumption $c_{i}<s$, utility level of an individual $i$ is negative, meaning that $i$’s life is not worth living.
-Hence, it must be that $c_{i}>s$ so the lifetime utility of all individuals are positive.
-Note however that under critical-level utilitarianism, an individual $i$’s life contributes positively to the value of a population if $\bar{c}<c_{i}$, while negatively if $s<c_{i}<\bar{c}$, even the person leads a life worth living.
-For analytical purposes, we employ ${\Upsilon}(c_{i})=\log c_{i}+\kappa$ with a $\kappa$ that the well-being subsistence $s=e^{-\kappa}$ is sufficiently close to zero.
-The population value of critical-level utilitarianism will then be $\mathcal{\mathcal{V}}=\sum_{i=1}^{n}\log c_{i}$ where $c_{i}\in[e^{-\kappa},\infty)$ and $\bar{c}=1$, which we hereafter seek to maximize, with respect to $c_{i}$ and $n$.
-As we recall our primary question, let us introduce descrete time in population ethics while assuming, for sake of simplicity, that periodical number of people is fixed to one unit; that is, all (representative) individuals are identified by the period $t$ they live in.
-[^2] We may therefore call $t$ a generation.
-The population ethics problem then becomes a generation ethics problem.
-The critical-level utilitarian population value function now becomes $\mathcal{\mathcal{V}}=\sum_{t=0}^{n}\log c_{t}$.
-We note that this type of value function (based on Bernoullian utility) is widely employed in the previous Ramsey-Cass-Koopmans type of dynastic optimization models (e.g., Nordhaus, 1992; Stern, 2007), except that the future generations are discounted according to the periodical deviation from the present, and typically, the number of potential generations $n$ is left unquestioned (and hence, set to infinity).
-In contrast, potential generations will be accounted equally whilst the number of generations will be subject to maximization, in this study.
+Economic models of intertemporal planning often assume an infinite horizon, implicitly treating dynasties or economies as perpetually forward-looking entities.
+However, such assumptions can obscure important trade-offs between current and future welfare, particularly when the viability of future generations depends on limited resources and capital dynamics.
+In this paper, we examine a social planner’s problem where the planning horizon—the number of generations considered—is endogenously chosen.
+By adopting a critical-level utilitarian welfare function, we evaluate the contribution of each generation relative to a minimum well-being threshold.
+This framework allows us to study how productivity, capital accumulation, and time preference interact to determine the ethically and economically optimal duration of a dynasty.
+We begin by positing a (classical) unitilitarian social planner whose objective is to indentify a state $x$ that maximizes the sum of the utilities $\mathcal{V}=\sum_{i\in n(x)}\Upsilon_{i}(x)$, which we refer to as the population value, where $n(x)$ denotes the set of individuals and $\Upsilon_{i}(x)$ denotes the lifetime utility (i.e., the level of well-being) of an individual $i$.
+For simplicity, let us assume that all individuals are identical and consume an equal share, denoted by $c$, from a fixed amount of resource $R>0$.
+A utilitarian axiology examines the problem of finding the population size $n=R/c$ that maximizes the population value $\mathcal{V}=\sum_{i=1}^{n}\Upsilon(c_{i})=n\Upsilon(c=R/n)$.
+The utility function $\Upsilon(c)$ is assumed to be monotonically increasing and concave.
+In addition, with respect to the population value $\mathcal{V}=n\Upsilon(c)$, zero utility ($\Upsilon=0$) is equivalent to the absence of life ($n=0$).
+The consumption level $c=\nu$, where $\Upsilon(\nu)=0$, be referred to as the well-being subsistence level, following Dasgupta (2019).
+Owing to the monotonicity of the utility function, $\Upsilon(c)>0$ must hold as long as $c>\nu$ (see Figure 1 (left)).
+It then follows that the domain allowed for $c$ is $(\nu,\infty)$ since $\Upsilon(c)$ represents the level of attainable lifetime utility.
+Figure 1 (left) depicts the optimal solution $c=\omega$, which satisfies the first-order condition for maximizing the population value, i.e., $\Upsilon^{\prime}(\omega)=\Upsilon(\omega)/\omega$.
+Owing to the concavity of the utility function, if the well-being subsistence approaches zero ($\nu\to 0$), as Córdoba (2023) refers to the homothetic case, the optimal consumption also approaches zero ($\omega\to\nu$), while the optimal size of the population approaches infinity ($n^{*}=R/\omega\to\infty$), leading to a conclusion that is considered repugnant (Parfit, 1986).
+Meanwhile, Blackorby et al (2005) introduced a population axiology known as critical-level utilitarianism.
+Within this framework, a social planner aims to determine a set of consumption schedules $c_{i}$ that maximizes the population value $\mathcal{V}=\sum_{i=1}^{n}\left(\Upsilon(c_{i})-\alpha\right)$, where an individual’s contribution to the population value is evaluated relative to the critical utility level $\alpha$ (see Figure 1 (right)).
+That is, an individual $i$’s life contributes positively to the population value if $c_{i}>\kappa=\Upsilon^{-1}(\alpha)$ but negatively otherwise.
+Note that setting $\alpha>0$ prevents the repugnant conclusion even when the underlying utility function is homothetic ($\nu=0$).
+For convenience, we employ a utility function of the form $\Upsilon(c_{i})=\log c_{i}+\alpha$ for analyses so that the value function for the critical-level utilitarian population value in our maximization study becomes $\mathcal{V}=\sum_{i=1}^{n}\left(\Upsilon(c_{i})-\alpha\right)=\sum_{i=1}^{n}\log c_{i}$.
+As we recall our primary idea of incorporating roundabout production into the optimal population framework, let us introduce (discrete) time and assume that the number of individuals per period is fixed to one unit; that is, all (representative) individuals are identified by the period $t$ during which they live.
+[^1] We may therefore refer to $t$ as a generation.
+The population value function then becomes $\mathcal{\mathcal{V}}=\sum_{t=0}^{n}\log c_{t}$.
+Note that this type of value function was widely employed in the previous Ramsey–Cass–Koopmans type of dynastic optimization model (including Nordhaus, 1992; Stern, 2007), except that the future generations were discounted according to the temporal deviation from the present, and typically, the number of potential generations $n$ remains unquestioned (and set to infinity).
+In contrast, in this study, potential generations are accounted for equally, and the number of generations is optimized.
 The remainder of the paper proceeds as follows.
-In the following section, we first specify the model of a production economy based on Cobb-Douglas technology from a dynamical perspective.
-We then specify our primary question in the form of a finite horizon dynamical programming problem where the objective is to maximize the sum of utilities given the abovementioned model of production, and derive a general solution function.
-In Section 3 we introduce two broad settings of parameters to simplify the general solution function.
-The first one is the AK setting where the output elasticity of capital is set to unity.
-The second one is the ZD (zero discounting) setting where the discount factor is set to unity.
-We also study the peculiarity of the two models in regard to intergenerational equity and optimal initial action of consumption.
+In the following section, we first describe the model of roundabout production on the basis of Cobb–Douglas technology.
+We then define our main problem in the form of finite horizon dynamic programming, where the objective function comprises a series of Bernoullian contributions, and analytically derive a general solution.
+In Section 3, we introduce two broad settings of parameters to simplify the general solution and allow us to study the optimal trajectory of the variables in a numerical fashion.
+The first setting is the AK setting, where the output elasticity of capital is set to unity.
+The second is the zero discounting (ZD) setting, where the discount factor is set to unity.
+We also study the peculiarity of the two models with respect to intergenerational equity and the optimal initial action of consumption.
 Section 4 concludes the paper.
-[…]
-Figure 1: Contribution to the population value (vertical axis) under critical-level utilitarianism.
-Left: The critical-level consumption $\bar{c}$ contributes none to the population value, i.e., ${\Upsilon}(\bar{c})-\kappa=0$.
-Consumption level $c^{*}$ maximizes the critical-level utilitarian population value.
-Right: Well-being subsistence $s$ bears zero lifetime utility (neutrality), i.e., ${\Upsilon}(s)=0$.
-2 The model
-2.1 The problem of a dynasity
-Let us begin by postulating a two-factor Cobb-Douglas production function of constant returns to scale, as follows:
-$Y_{t}={A}(K_{t})^{\theta}(L_{t})^{1-\theta}=\frac{{B}}{\theta^{\theta}(1-\theta)^{1-\theta}}(K_{t})^{\theta}(L_{t})^{1-\theta}$ (1)
-where $Y_{t}$ denotes total output of the economy, $K_{t}$ denotes capital, and $L_{t}$ denotes labor, all effective in period $t$.
-As for the parameters concerned, $\theta<1$ denotes the output elasticity of capital, ${A}>0$ denotes the level of technology termed productivity, and ${B}={A}\Theta$, where $\Theta=\theta^{\theta}(1-\theta)^{1-\theta}$, denotes the cost function-based productivity, which we otherwise call the dual productivity.333 See Appendix 1 for more details.
-Below describes the breakdown of total output into investment, depreciation of capital, and final consumption:
+Figure 1: Left: Utility function of an individual.
+The consumption level for neutrality (or well-being subsistence) is denoted by $\nu$.
+Marginal utility and average utility coincide at $c=\omega$.
+Right: Contribution to the critical-level utilitarian population value, where $\alpha$ denotes the critical level.
+At $c=\kappa$, the utility reaches the critical level.
+In all cases, the domain of permissible consumption levels $c$ for existence is $c>\nu$.
+2 Model
+2.1 Roundabout production
+Let us begin by postulating a two-factor Cobb–Douglas production function characterized by constant returns to scale, as follows:
+$Y_{t}={A}(K_{t})^{\theta}(L_{t})^{1-\theta}$ (1)
+where $Y_{t}$ denotes the economy’s output, $K_{t}$ denotes capital, and $L_{t}$ denotes labor, all of which are effective during period $t$.
+For the relevant parameters, $10<\theta<$ 1 denotes the output elasticity of capital.
+The level of technology is indicated by productivity, denoted by $A$, which is assumed to be fixed throughout the considered time span.
+Note that ${B}\equiv{A}\theta^{\theta}(1-\theta)^{1-\theta}$ is the cost function-based productivity.
+[^2] Otherwise, $B$ is referred to as the dual productivity.
+See Appendix 1 for more details.
+The breakdown of output into investment, capital depreciation, and final consumption is described below:
 $Y_{t}=K_{t+1}-K_{t}+\delta K_{t}+C_{t}$ (2)
-Here, $\delta\leq 1$ denotes depreciation factor of capital, and $C_{t}$ denotes consumption in period $t$.
-Combining equations (1) and (2), and dividing both sides by $L_{t}=L$ which we assume to be constant over time, leads to the following intertemporal dynamics of capital intensity:
+Here, $\delta\leq 1$ denotes the capital depreciation factor, and $C_{t}$ denotes aggregate consumption in period $t$.
+Combining equations (1) and (2) and dividing both sides by $L_{t}=L$, which we assume to be constant over time, leads to the following intertemporal dynamics (or roundaboutness) of capital intensity:
 $y_{t}={A}(k_{t})^{\theta}=k_{t+1}-(1-\delta)k_{t}+c_{t}$ (3)
-where $k_{t}=K_{t}/L$ and $k_{t+1}=K_{t+1}/L$ denote capital intensities in period $t$ and $t+1$, respectively.
-Also, $y_{t}=Y_{t}/L$ and $c_{t}=C_{t}/L$ denote per caput output and consumption, respectively, in period $t$.
+where $k_{t}=K_{t}/L$ and $k_{t+1}=K_{t+1}/L$ denote capital intensities in periods $t$ and $t+1$, respectively.
+Additionally, $y_{t}=Y_{t}/L$ and $c_{t}=C_{t}/L$ denote per capita output and consumption, respectively, in period $t$.
 […]
 2.2 Finite horizon dynamic programming
-Our primary problem here is to solve for an optimal consumption path given any planning horizon.
-The Bellman equation of the problem is of the following:555 Note that the objective function of the primary problem (2.1) is given at $t=0$, i.e., $\mathcal{V}[n]=\mathcal{V}_{0}[k_{0};n]$.
+The primary problem here is to solve for an optimal consumption path given a planning horizon.
+The Bellman equation of the problem is as follows:444 Note that the objective function of the primary problem (2.1) is given at $t=0$, i.e., $\mathcal{V}[n]=\mathcal{V}_{0}[k_{0};n]$.
 $\mathcal{\mathcal{V}}_{t}\left[k_{t};\ n\right]=\max_{c_{t}}\left(\log c_{t}+\beta\mathcal{\mathcal{V}}_{t+1}\left[k_{t+1}={A}(k_{t})^{\theta}-c_{t};\ n\right]\right)$ (4)
-The optimal trajectory of the state variable $k_{t}$ is given, by Lemma 2 which we append to Appendix 2 with proofs, as follows:
+The optimal trajectory of the state variable $k_{t}$ is given by Lemma 2, which we append to Appendix 2 with proofs, as follows:
 $k_{t}^{*}[n]=(k_{0})^{\theta^{t}}\prod_{i=1}^{t}\left(\frac{S_{{n}-i}}{S_{{n}-i+1}}{A}\beta\theta\right)^{\theta^{t-i}}=(k_{0})^{\theta^{t}}\prod_{i=1}^{t}\left(\frac{S_{{n}-t+i-1}}{S_{{n}-t+i}}{A}\beta\theta\right)^{\theta^{i-1}}$ (5)
@@
 $S_{\ell}\equiv\sum_{i=0}^{\ell}(\beta\theta)^{i}=1+(\beta\theta)+(\beta\theta)^{2}+\cdots+(\beta\theta)^{\ell}=\frac{1-(\beta\theta)^{\ell+1}}{1-\beta\theta}$ (6)
-The following optimal trajectory of consumption is obtained by (5), and (21) which must be true by the proof of Lemma 1.
+The following optimal trajectory of consumption is obtained by (5) and (21) Appendix 2, which must be true according to the proof of Lemma 1.
 $c_{t}^{*}[n]=\frac{{A}\left(k_{t}^{*}[n]\right)^{\theta}}{S_{n-1}}=\frac{{A}(k_{0})^{\theta^{t+1}}}{S_{{n}-t}}\prod_{i=1}^{t}\left(\frac{S_{{n}-t+i-1}}{S_{{n}-t+i}}{A}\beta\theta\right)^{\theta^{i}}$ (7)
-With formula (7) the population value function is given as follows:
+With formula (7), the population value function is given as follows:
 $\mathcal{V}[n]$ $=\sum_{t=0}^{n}\beta^{t}\log c_{t}^{*}[n]$
 $=\log\left(\frac{{A}(k_{0})^{\theta}}{S_{n}}\right)+\sum_{t=1}^{n}\beta^{t}\log\left(\frac{{A}(k_{0})^{\theta^{t+1}}}{S_{{n}-t}}\prod_{i=1}^{t}\left(\frac{S_{{n}-t+i-1}}{S_{{n}-t+i}}{A}\beta\theta\right)^{\theta^{i}}\right)$ (8)
-We hereafter seek to maximize this function with respect to the planning horizon $n$.
-Our approach to the analysis of population value (8) will be numerical rather than analytical, because the derivative for $\mathcal{V}[n]$ with respect to $n$ seems uninformative.
-As it turns out in the following section, the optimal tragectory of contributions $\log c_{t}^{*}[n]$ and the population value $\mathcal{V}[n]$, for any given planning horizon $n$, becomes manageable under $\theta=1$ (known as the AK setting) and $\beta<1$.
-We also find that the trajectory of contributions $\log c_{t}^{*}[\infty]$ and the population value $\mathcal{V}[\infty]$ for an infinite planning horizon, is evaluable under $\beta\theta<1$.
-We therefore base our study of zero discounting ($\beta=1$) on this setting (i.e., $\beta\theta<1$, which indicates $\theta<1$, a Cobb-Douglas model).
-In the following section, we delve into the abovementioned two broad settings of parameters, namely, AK production with future discounting ($\theta=1$ and $\beta<1$) which we term as AK settings, and Cobb-Douglas production without future discounting ($\theta<1$ and $\beta=1$) which we term as ZD settings.
-Table 1 summarizes the parameter settings chosen for numerical examinations.
-Note that cases I–IV correspond to AK settings whose solution paths are characterized by the sign of $\log({A}\beta)$, whereas cases V–VII correspond to ZD settings whose solution paths are characterized by the sign of $\log(A\Theta)$.
-The final case VIII corresponds to the parameter settings that exemplifies $\log({A}\beta)=\log(A\Theta)=0$.
+We hereafter aim to maximize this function with respect to the planning horizon $n$.
+Our approach to analyzing the population value (8) adopts a numerical rather than an analytical framework because the derivative of $\mathcal{V}[n]$ with respect to $n$ does not seem to provide meaningful insights.
+In the following section, the optimal trajectory of contributions $\log c_{t}^{*}[n]$ and the population value $\mathcal{V}[n]$, for any given planning horizon $n$, becomes manageable under $\theta=1$ (known as the AK setting) and $\beta<1$.
+We also find that the trajectory of contributions $\log c_{t}^{*}[\infty]$ and the population value $\mathcal{V}[\infty]$ for an infinite planning horizon is evaluable under $\beta\theta<1$.
+We therefore base our study of ZD ($\beta=1$) on this parameter setting (i.e., $\beta\theta<1$, which indicates that $\theta<1$, consistent with a Cobb–Douglas model).
+In the following section, we delve into the abovementioned two broad settings of parameters, namely, AK production with future discounting ($\theta=1$ and $\beta<1$), which we term AK settings, and Cobb–Douglas production without future discounting ($\theta<1$ and $\beta=1$), which we term ZD settings.
+Table 1 summarizes the parameter settings chosen for the numerical examinations.
+Note that cases I–IV correspond to AK settings whose solution paths are characterized by the sign of $\log(A\beta)$, whereas cases V–VII correspond to ZD settings whose solution paths are characterized by the sign of $\log B$.
+Cases VIII and IX correspond to the parameter settings where $\log({A}\beta)=\log B=0$.
 Table 1: Parameters applied in various cases.
-case ${A}$ $\beta$ $\theta$ $\log({A}\beta)$ $\log({A\Theta})$ $n^{*}$ $\mathcal{V}[n^{*}]$ $\mathcal{V}[\infty]$
-I $1.012$ $0.992$ $1$ $+$ $\infty$ $84.675$
-II $1.01$ $0.992$ $1$ $+$ 95 $59.965$
-III *1 $0.992$ $1$ 00 73 $55.627$
-IV $1.005$ $0.992$ $1$- 58 $50.979$
+case ${A}$ $\beta$ $\theta$ $\log({A}\beta)$ $\log B$ $n^{*}$ $\mathcal{V}[n^{*}]$ $\mathcal{V}[\infty]$
+I $1.012$ $0.992$ $1$ $+$ $\infty$ $84.7$
+II $1.01$ $0.992$ $1$ $+$ $95$ $60.0$
+III $1/\beta$ $0.992$ $1$ 00 $73$ $55.6$
+IV $1.005$ $0.992$ $1$- $58$ $51.0$
 V $1.05$ $1$ $0.992$ $+$ $\infty$ +$\infty$+ ∞
-VI $1.2$ $1$2 00 (281) $41.688$
+VI $1.2$ $1$1 00 $(281)$ $41.7$
 VII $1.05$ $1$ $0.991$- $117$$\infty$- ∞
-VIII $1$ $1$ $1$ 00 00 54 $55.182$$\infty$- ∞
+VIII $1$ $1$ $1$ 00 00 $54$ $55.2$$\infty$- ∞
+IX $1/\beta$ $0.992$2 00 00 $53$ $49.1$ $-1843.2$
   • Note:
-    For all cases I–VIII, $k_{0}=150$.
+    For all cases I–IX, $k_{0}=150$. $B=A(1-\theta)^{1-\theta}\theta^{\theta}$ is the cost-function based productivity.
   • *1
-    ${A}=1/\beta\approx 1.008065$.
+    $\theta\approx 0.955392$ where 1.2($1{1.2}(1-\theta)^{1-\theta}\theta^{\theta}=B=$ 1.
   • *2
-    $\theta\approx 0.955392$ where 1.2($1{1.2}(1-\theta)^{1-\theta}\theta^{\theta}=$ 1.
+    $\theta\approx 0.998982$ where $(1/0.992)(1-\theta)^{1-\theta}\theta^{\theta}=B=1$.
 3 Analysis
 3.1 AK setting
-AK model is one of the simplest yet fundamental models of endogenous growth, which was formally developed by Frankel (1962).
-Here, we take this production model to study the population value function that discounts future generations to see whether if $n\to\infty$ is an optimal policy.
-The optimal consumption path (7) for AK setting with future discounting ($\theta=1$, $\beta<1$) may be specified as follows:
+The AK model, which was formally developed by Frankel (1962), is one of the simplest fundamental models of endogenous growth.
+Here, we employ this production model to study the population value function that discounts future generations to determine whether $n\to\infty$ is an optimal policy.
+The optimal consumption path (7) for the AK setting with future discounting ($\theta=1$, $\beta<1$) may be specified as follows:
 $c_{t}^{*}[n]=\left(\frac{{A}k_{0}}{S_{{n}-t}}\right)\left(\frac{S_{{n}-1}}{S_{n}}{A}\beta\right)\left(\frac{S_{{n}-2}}{S_{{n}-1}}{A}\beta\right)\cdots\left(\frac{S_{{n}-t}}{S_{{n}-t+1}}{A}\beta\right)=\frac{({A}\beta)^{t}{A}k_{0}}{S_{n}}$ (9)
@@
 $=\frac{\beta-((1-\beta){n}+1)\beta^{{n}+1}}{(1-\beta)^{2}}\log({A}\beta)+\frac{1-\beta^{{n}+1}}{1-\beta}\log\left(\frac{1-\beta}{1-\beta^{{n}+1}}{A}k_{0}\right)$ (10)
-For sake of the analysis let us take the derivative with respect to ${n}$.
+For the sake of the analysis, let us take the derivative with respect to ${n}$.
 $\frac{\mathrm{d}\mathcal{V}[n]}{\mathrm{d}n}=\lambda\left(\gamma-{n}\log({A}\beta)+\log\left(1-\beta^{n+1}\right)\right)$ (11)
-where $\lambda$ and $\gamma$, that are both independent of $n$, are given as follows:
+where $\lambda$ and $\gamma$ are given as follows:
 $\lambda=\frac{\beta^{n}\log\beta^{\beta}}{1-\beta}>0,$ $\gamma=1-\frac{\log({A}\beta)^{1-\beta+\log\beta}}{\log\beta^{1-\beta}}-\log({A}(1-\beta)k_{0})$
-In order to study the derivative sign of (11), we consider the following two functions whose equality convey the first order conditions, i.e., $f[n]=g[n]$.
+To study the derivative sign of (11), we consider the following two functions, whose equal values, i.e., $f[n]=g[n]$, convey the first-order condition of optimality.
 $f[n]=-\gamma+n\log({A}\beta),$ $g[n]=\log\left(1-\beta^{n+1}\right)$ (12)
 Before we proceed, let us examine the possible range of the parameter $A$.
-Described below is the evaluation of the marginal product of capital (MPK) of our production function (1) net of depreciation, which should coincide with the real interest rate $\rho>0$, viz.,
+The evaluation of the marginal product of capital (MPK) of our production function (1) net of depreciation, which should coincide with the real interest rate $\rho>0$, viz.,
 $\frac{\partial Y_{t}}{\partial K_{t}}-\delta={A}\theta\left(\frac{K_{t}}{L_{t}}\right)^{\theta-1}-\delta=\rho>0$ (13)
-Then, we know that AK models with complete capital depreciation $\theta=\delta=1$ amounts to ${A}=1+\rho$, and therefore ${A}>1$ is a relevant assumption in this setting.666 Note also that $A\beta=(\delta+\rho)\beta$ indicates the discrepancy between the gross rate of interest and the rate of generational preference of the population.
-Figure 2 (left) depicts functions $f[n]$ and $g[n]$ under different parameters for cases I–IV.
-Clearly, $f[n]$ is a linear function whose slope is $\log({A}\beta)$, and $g[n]$ is monotonously increasing and approaching zero, i.e., $g[\infty]=0$.
-Here, we fix the discount factor $\beta$, and differentiate the productivity ${A}$ in four different levels.
-As long as ${A}\leq 1/\beta$, so that $\log({A}\beta)\leq 0$, i.e., the slope is zero or negative, $f[n]$ will intersect with $g[n]$ at a single point and the population value function $\mathcal{V}[n]$ has a single peak.
-If ${A}>1/\beta$ so that $f[n]$ has a positive slope, then $f[n]$ and $g[n]$ could either intersect with two points where the population value $\mathcal{V}[n]$ would rise, fall and then rise again, or never intersect so that the population value would be rising indefinitely, with respect to the planning horizon $n$.
+In AK models with complete capital depreciation, $\theta=\delta=1$ leads to ${A}=1+\rho$.
+Therefore, assuming that ${A}>1$ is relevant in this setting.555 Note also that $A\beta=(\delta+\rho)\beta$ indicates the discrepancy between the gross interest rate and generational preference rate of the population.
+Figure 2 (left) depicts the functions $f[n]$ and $g[n]$ under different parameters for cases I–IV.
+Clearly, $f[n]$ is a linear function whose slope is $\log({A}\beta)$, and $g[n]$ monotonously increases and approaches zero, i.e., $g[\infty]=0$.
+Here, we fix the discount factor $\beta$ and differentiate the productivity ${A}$ at four different levels.
+As long as ${A}\leq 1/\beta$, so that $\log({A}\beta)\leq 0$, i.e., the slope is zero or negative, $f[n]$ will intersect with $g[n]$ at a single point, and the population value function $\mathcal{V}[n]$ has a single peak.
+If ${A}>1/\beta$ so that $f[n]$ has a positive slope, then $f[n]$ and $g[n]$ could either intersect with two points where the population value $\mathcal{V}[n]$ may rise, fall and then rise again, or never intersect so that the population value may rise indefinitely with respect to the planning horizon $n$.
 Figure 2 (right) depicts the population value functions for cases I–IV.
-Figure 2: Left: Solid plots show $g[n]$ defined in (12) for cases I–IV with obvious correspondences, i.e., the steepest one corresponds to case I, whereas case IV corresponds to the one with a negative slope.
-Dashed plot shows $f[n]$ also defined in (12).
-Right: The plots show population value paths $\mathcal{V}[n]$ for cases I–IV with obvious correspondences.
+Figure 2: Left: The solid lines represent $g[n]$, as defined in (12), for cases I–IV with clear correspondences, i.e., the steepest slope corresponds to case I, while case IV corresponds to the scenario with a negative slope.
+The dashed line represents $f[n]$, which is also defined in (12).
+Right: The plots show population value functions $\mathcal{V}[n]$ for cases I–IV with obvious correspondences.
 To visualize the optimal trajectory of the variables in various situations, we specify them here in the form of functionals.
-By virtue of (9), the optimal trajectory of undiscounted contribution to the population value becomes linear with respect to $t$, as follows:
+By referencing (9), the optimal trajectory of the undiscounted contribution to the population value becomes linear with respect to $t$, as follows:
 $\log c_{t}^{*}[n]=\log\left(\frac{1-\beta}{1-\beta^{n+1}}{A}k_{0}\right)+t\log({A}\beta)$ (14)
-By virtue of (21), the optimal trajectory of capital intensity becomes as follows:
+In reference to (21) Appendix 2, the optimal trajectory of capital intensity becomes as follows:
 $k_{t}^{*}[n]=\frac{S_{n-t}}{{A}}c_{t}^{*}[n]=\frac{1-\beta^{n+1-t}}{1-\beta^{n+1}}({A}\beta)^{t}k_{0}$ (15)
-Figure 3 displays the population value function $\mathcal{V}[n]$, specified as (10), for cases I, III, and IV, on the top.
-[^7] These figures are reiterations from Figure 2 (right).
-Note that, as long as $\beta<1$, the population value function will always converge to some finite value, regardless of the parameter settings besides $\beta$.
-As we let $n\to\infty$ in (10), we have what follows:
+Figure 3 displays the population value function $\mathcal{V}[n]$, specified as (10), for cases I, III, and IV on the top.
+[^6] These figures correspond to those depicted in Figure 2 (right).
+Note that, as long as $\beta<1$, the population value function will always converge to a finite value, regardless of the parameter settings in addition to $\beta$.
+As we let $n\to\infty$ in (10), we have the following:
 $\mathcal{V}[\infty]=\frac{\log\left({A}(1-\beta)^{1-\beta}\beta^{\beta}\right)}{(1-\beta)^{2}}+\frac{\log(k_{0})}{1-\beta}$ (16)
-By comparing the population values at the numerical solution of the first order condition $f[n]=g[n]$ in reference to (12), with $\mathcal{V}[\infty]$ in reference to (16), we know the optimal planning horizon $n^{*}$, which we display in Table 1 for cases I–IV.
-At the middle row of Figure 3, optimal trajectories of undiscounted contribution to the population value, according to (14), with respect to planning horizons $n=200,400,600$, are displayed for sample I, III, and IV from left to right.
-Similarly at the bottom row of Figure 3, optimal trajectories of capital intensity, according to (15), with respect to planning horizons $n=200,400,600$, are displayed for cases I, III, and IV from left to right.
-Figure 3: Population value function (top), trajectories of undiscounted contribution (middle) and capital intensity (bottom), for AK model cases I (left), III (center), and IV (right).
+By comparing the population values obtained at the numerical solution of the first-order condition $f[n]=g[n]$, as referenced in (12), with $\mathcal{V}[\infty]$ from (16), we know the optimal planning horizon $n^{*}$, which we display in Table 1 for cases I–IV.
+The middle row of Figure 3 displays the optimal trajectories of the undiscounted contributions to the population value, as described in (14), for planning horizons $n=200,400,600$ for cases I, III, and IV (from left to right).
+Similarly, the bottom row of Figure 3 displays the optimal trajectories of capital intensity, based on (15), for planning horizons $n=200,400,600$ for cases I, III, and IV (from left to right).
+Figure 3: Population value function (top), trajectories of undiscounted contribution (middle) and capital intensity (bottom), for AK setting cases I (left), III (center), and IV (right).
+The parameters corresponding to each of the cases are given in Table 1.
 3.2 ZD setting
-Here, we study zero discounting $\beta=1$ under Cobb-Douglas production $\theta<1$. The condition that $\beta\theta<1$ nonetheless allows us to evaluate the key summation as follows:
+Here, we study ZD, i.e., $\beta=1$, under Cobb–Douglas production with an ordinary output elasticity of capital $\theta<1$.
+The condition that $\beta\theta<1$ nonetheless allows us to evaluate the key summation as follows:
 $S_{\infty-\tau}=\lim_{n\to\infty}\frac{1-(\beta\theta)^{n-\tau+1}}{1-\beta\theta}=\frac{1}{1-\beta\theta}$
-As we apply the above and $n\to\infty$ into (7), we have the following:
+We apply the above and $n\to\infty$ to (7) and obtain the following:
 $\log c_{t}^{*}[\infty]$ $=\log\left(\frac{{A}}{S_{{\infty}-t}}\right)+\theta\left(\sum_{i=1}^{t}\theta^{t-i}\log\left(\frac{S_{{\infty}-i}}{S_{{\infty}-i+1}}{A}\beta\theta\right)+\theta^{t}\log k_{0}\right)$
@@
 $=\frac{\log\left({A}(1-\beta\theta)^{1-\theta}(\beta\theta)^{\theta}\right)}{1-\theta}+\theta^{t}\left(\log(k_{0})^{\theta}-\log({A}\beta\theta)^{\frac{\theta}{1-\theta}}\right)$ (17)
-where the optimal trajectory for undiscounted contributions turns out to be geometrically convergent.
-We can then apply $\beta=1$ to arrive at the following zero discount verstion of the optimal trajectory for undiscounted contributions:
+where the optimal trajectory for undiscounted contributions is geometrically convergent.
+We can then apply $\beta=1$ to arrive at the following zero discount verification of the optimal trajectory for undiscounted contributions:
 $\log c_{t}^{*}[\infty]=\frac{\log\left({A}(1-\theta)^{1-\theta}\theta^{\theta}\right)}{1-\theta}+\theta^{t}\left(\log(k_{0})^{\theta}-\log({A}\theta)^{\frac{\theta}{1-\theta}}\right)$
@@
 $\mathcal{V}[\infty]=\sum_{t=0}^{\infty}\log c_{t}^{*}[\infty]=$ $\frac{\infty\log\left({A}(1-\theta)^{1-\theta}\theta^{\theta}\right)+\log(k_{0})^{\theta}-\log({A}\theta)^{\frac{\theta}{1-\theta}}}{1-\theta}$
-$=$ $\begin{cases}+\infty&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}={A}\Theta>1\\ {\frac{\theta}{1-\theta}}\log\left(\frac{1-\theta}{\theta}k_{0}\right)&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}={A}\Theta=1\\ -\infty&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}={A}\Theta<1\end{cases}$
-The above result indicates that if ${A}\Theta>1$, then the population value is ever increasing and $n\to\infty$ must be the optimal solution; whereas, if ${A}\Theta<1$, then $n\to\infty$ must not be optimal and the population value must be maximized at a finite horizon $n^{*}\ll\infty$.
-The case where ${A}\Theta=1$ turns out to be the one that the planning horizon $n$ does not matter (beyond certain length) in maximizing the population value.
-Case V reflects ${A}\Theta>1$ where $n^{*}=\infty$, case VII reflects ${A}\Theta<1$ where $n^{*}\ll\infty$, and case VI reflects ${A}\Theta=1$ where $n^{*}$ is any number greater than $281$.888 The number is subject to decimals rounding.
-Figure 4 dipicts the population value function $\mathcal{V}[n]$, the optimum trajectory of undiscounted contribution $\log c_{t}^{*}[n]$ and that of capital intensity $k_{t}^{*}[n]$, given the planning horizon $n$, for sample ZD settings V, VI, and VII.
-The final case VIII implies ZD and AK setting ($\theta=\beta=1$), where $A\Theta=A=1+\rho>1$, under complete depreciation, by virtue of (13).
-Thus, if $\rho>0$ then $A>1$ in which case the population value would explode as described above, and we are left with $n^{*}\to\infty$.
-If in turn, $A=1+\rho=1$ as the case (VIII) goes, the infinite horizon population value is evaluated as follows:
+$=$ $\begin{cases}+\infty&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}=B>1\\ {\frac{\theta}{1-\theta}}\log\left(\frac{1-\theta}{\theta}k_{0}\right)&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}=B=1\\ -\infty&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}=B<1\end{cases}$
+The above result indicates that if $B>1$, then the population value is ever increasing, and $n\to\infty$ must be the optimal solution.
+However, if $B<1$, then $n\to\infty$ must not be optimal, and the population value must be maximized at a finite horizon $n^{*}\ll\infty$.
+The case where $B=1$ is the knife-edge case in which the planning horizon $n$ does not matter (beyond a certain length) in maximizing the population value.
+Case V corresponds to $B>1$, and $n^{*}=\infty$; case VII corresponds to $B<1$, and $n^{*}\ll\infty$; and case VI corresponds to $B=1$, and $n^{*}$ is any number greater than $281$.777 The number is subject to decimal rounding.
+Figure 4 depicts the population value function $\mathcal{V}[n]$, along with the optimum trajectory of undiscounted contribution $\log c_{t}^{*}[n]$ and the capital intensity $k_{t}^{*}[n]$, given a planning horizon $n$, for sample ZD settings V, VI, and VII.
+The final case VIII relates to the ZD/AK setting ($\theta=\beta=1$), where $B=A=1+\rho>1$, under complete depreciation, as described in (13).
+Thus, if $\rho>0$, then $A>1$, in which case the population value would increase without bound, as described above, leading to $n^{*}\to\infty$.
+If, in turn, $A=1+\rho=1$, as in case VIII, the infinite horizon population value is evaluated as follows:
 $\mathcal{V}[\infty]=\lim_{\theta\to 1}{\frac{\theta}{1-\theta}}\log\left(\frac{1-\theta}{\theta}k_{0}\right)=-\infty$
 This indicates that there exists a finite optimum horizon $n^{*}\ll\infty$.
-Note that the population value function in this setting can be specified by zero discounting $(\beta\to 1)$ the AK model’s population value function (10), as follows:
+Note that the population value function in this setting can be specified by applying a unitary discounting factor $(\beta\to 1)$ to the population value function (10) of AK models as follows:
 $\mathcal{V}[n]$ $=\lim_{\beta\to 1}\left(\frac{\beta-((1-\beta){n}+1)\beta^{{n}+1}}{(1-\beta)^{2}}\log({A}\beta)+\frac{1-\beta^{{n}+1}}{1-\beta}\log\left(\frac{1-\beta}{1-\beta^{{n}+1}}{A}k_{0}\right)\right)$
@@
 The third identity is subject to L’Hôpital’s rule.
-By the first order condition $\frac{\partial\mathcal{V}[n]}{\partial n}=0$, the optimal planning horizon is evaluated as $n^{*}=\exp(-1+\log k_{0})-1\approx 54.182$.
+By the first-order condition $\frac{\partial\mathcal{V}[n]}{\partial n}=0$, the optimal planning horizon is evaluated as $n^{*}=\exp(-1+\log k_{0})-1\approx 54.182$.
 Figure 4: Population value function (top), trajectories of undiscounted contribution (middle) and capital intensity (bottom), for ZD setting cases V (left), VI (center), and VII (right).
 Note that the population value function of the top-left panel is increasing indefinitely.
+The parameters corresponding to each of the cases are given in Table 1.
 3.3 Intergenerational inequality
-It seems obvious from Figure 3 (middle row) that consumption inequality among generations is worsened as we expand the planning horizon in the presence of future discounting (except for the knife-edge case III with $\log(A\beta)=0$), while it seems to improve in Figure 4 (middle row), for cases with zero discounting.
-To check the validity of the conjecture, we prepare Figure 5 to illustrate how Lorenz curve moves in regard to the planning horizon (i.e., $n=200,400,600$) for AK setting cases I, III, IV, and for ZD setting cases V, VI, and VII.
-The Lorenz curves (in Figure 5) are based on the stream of optimal consumption for a given planning horizon, namely, $c_{t}^{*}[n]$.
-The level of inequality increases in AK settings as the planning horizon is stretched longer.
-On the other hand, the level of inequality is relatively insensitive to the planning horizon for zero discounting cases, except for case VI with $\log(A\Theta)=0$ where inequality decreases along with the length of the planning horizon.
+Figure 3 (middle row) shows that consumption inequality across generations increases as the planning horizon extends under future discounting (except in the knife-edge case III when $\log(A\beta)=0$), whereas it seems to decrease in Figure 4 (middle row) for cases with ZD.
+To validate this conjecture, Figure 5 shows how the Lorenz curve shifts with respect to the planning horizon (namely, $n=200,400,600$) for AK setting cases I, III, and IV and ZD setting cases V, VI, and VII.
+These Lorenz curves (of Figure 5) are based on the sequence of optimal consumption levels for a given planning horizon, i.e., $c_{t}^{*}[n]$.
+The inequality level increases in AK settings as the planning horizon expands.
+On the other hand, inequality level is relatively insensitive to the planning horizon for ZD setting cases, except for case VI, when $\log B=0$, where inequality decreases as the planning horizon increases.
 Figure 5: Lorenz curves for $n=200$ (solid line), $n=400$ (dotted line), and $n=600$ (dashed line).
-Top row panels (from left to right) correspond to AK settings I, III, and IV.
-Second row panels (from left to right) correspond to ZD settings V, VI, and VII.
-For further analysis, let us consider Gini Index, a popular measure of inequality, defined on the optimal stream of consumptions $c_{t}^{*}[n]$, as follows:
+The top row panels (from left to right) correspond to AK settings I, III, and IV.
+The second row panels (from left to right) correspond to ZD settings V, VI, and VII.
+For further analysis, let us consider the Gini index, a popular measure of inequality, defined on the basis of the optimal stream of consumption $c_{t}^{*}[n]$, as follows:
 $\mathcal{G}[n]=\frac{\sum_{t^{\prime}=0}^{n}\sum_{t=0}^{n}\left|c_{t}^{*}[n]-c_{t^{\prime}}^{*}[n]\right|}{2n\sum_{t=0}^{n}c_{t}^{*}[n]}$
-Figure 6 illustrates Gini Index $\mathcal{G}$ as a function of planning horizon $n$ in various parameter settings.
-For both figures the the underlying parameters are fixed at $k_{0}=150$ and $A=1$.
-The left figure pertains to AK settings ($\theta=1$), while we raise the discount rate from $\beta=0.9$ (solid line) to $\beta=0.99$ (dashed line).
-As we can see from the (left) panel, AK settings have the tendency to unequalize consumptions between generations as the planning horizon expands, while inequality is mitigated by less discountig the future.
-In contrast, the right figure indicates that ZD settings have the tendency to equalize consumptions between generations as the planning horizon expands, while inequality is enhanced by the higher output elasticity of capital.
-Figure 6: The panels are planning horizon vs Gini Index i.e., $(n,\mathcal{G}[n])$ for AK model cases with different discount rates (left), and zero discounting cases with different output elasticities of capital (right).
-The parameters setting of the left panel is $\theta=1$, $k_{0}=150$, $A=1$, $\beta=0.9$ (solid line), and $\beta=0.99$ (dashed line).
-The parameters setting of the right panel is $\beta=1$, $k_{0}=150$, $A=1$, $\theta=0.9$ (solid line), and $\theta=0.95$ (dashed line).
-3.4 The initial action
-Let us now focus on the initial action of the optimal consumption schedule which we specify, by virtue of (7) or (21), as follows:
+Figure 6 shows the Gini index $\mathcal{G}$ as a function of the planning horizon $n$ under various parameter settings.
+For both panels, the underlying parameters are fixed at $k_{0}=150$ and $A=1$.
+The left panel corresponds to the AK settings ($\theta=1$) as the discount rate is increased from $\beta=0.9$ (solid line) to $\beta=0.99$ (dashed line).
+In the AK settings, consumption tends to become more unequal across generations as the planning horizon expands.
+However, this effect is mitigated when future discounting decreases.
+In contrast, the right panel indicates that ZD settings tend to equalize consumption across generations as the planning horizon expands, whereas inequality is enhanced by the higher output elasticity of capital.
+Figure 6: The panels depict the planning horizon vs Gini Index, i.e., $(n,\mathcal{G}[n])$ for AK setting cases with different discount rates (left), and ZD setting cases with different output elasticities of capital (right).
+The parameter settings for the left panel are $\theta=1$, $k_{0}=150$, $A=1$, $\beta=0.9$ (solid line), and $\beta=0.99$ (dashed line).
+The parameter settings for the right panel are $\beta=1$, $k_{0}=150$, $A=1$, $\theta=0.9$ (solid line), and $\theta=0.95$ (dashed line).
+3.4 Initial action
+Let us focus on the initial action of the optimal consumption schedule, which we specify as follows based on (7) or (21):
 $c_{0}^{*}[n]=\frac{{A}(k_{0})^{\theta}}{S_{n}}=\frac{1-\beta\theta}{1-(\beta\theta)^{n}}{A}(k_{0})^{\theta}$ (18)
@@
 $\frac{\mathrm{d}c_{0}^{*}[n]}{\mathrm{d}n}=\frac{A(k_{0})^{\theta}\left(1-\beta\theta\right)(\beta\theta)^{n}\log(\beta\theta)}{\left(1-(\beta\theta)^{n}\right)^{2}}<0$
-since, $\beta\theta<1$.
-In other words, the further we take future generations into account, the more current generation must sacrifice.
-In any case, the initial action is monotonously decreasing in the planning horizon, and ultimately converge to the following:
+since $\beta\theta<1$.
+In other words, the more future generations are treated equally, the more current generation must reduce their consumption.
+In any case, the initial action monotonically decreases in the planning horizon and ultimately converges to the following:
 $c_{0}^{*}[\infty]=(1-\beta\theta){A}(k_{0})^{\theta}$ (19)
-By taking it’s derivative with respect to $\beta$ we obtain:
+By taking its derivative with respect to $\beta$, we have:
 $\frac{\partial c_{0}^{*}[\infty]}{\partial\beta}=-\theta A(k_{0})^{\theta}<0$
-That is to say, the more we respect future generations (by raising $\beta$), the more current generation must sacrefice.
-From another perspective, we can solve the equation $c_{0}^{*}[n]=s$ for $n$ in regard to (18), where $s$ is the consumption level of subsistence, as follows:
-$n=\frac{\log\left((1-\beta\theta)A(k_{0})^{\theta}\right)-\log s}{\log(\beta\theta)}$
-This $n$ is the subsistence-proof number of potential generations.
-Finally, for AK models, $c_{0}^{*}[\infty]=(1-\beta)Ak_{0}$ by (19), but this value will approach zero if future generations are respected as equally as possible to the current one ($\beta\to 1$).
-That is to say, the ultimate respect for the future generations in AK models with an infinite planning horizon induces the current generation to sacrefice ultimately, i.e., $c_{0}^{*}[\infty]\to{s}$.
-Alternatively, if we relax the infinite horizon assumption, initial action for AK models with ultimate respect for the future generations be:
+In other words, the more we consider future generations (by raising $\beta$), the more the current generation must reduce their consumption.
+From another perspective, we can solve equation $c_{0}^{*}[n]=\nu$ for $n$ with respect to (18), where $\nu$ is the well-being subsistence, as follows:
+$n=\frac{\log\left((1-\beta\theta)A(k_{0})^{\theta}\right)-\log\nu}{\log(\beta\theta)}$
+This $n$ is the subsistence-proof size of potential generations.
+Finally, for AK models, the optimal initial action by (19) is $c_{0}^{*}[\infty]=(1-\beta)Ak_{0}$, but this value approaches zero if future generations are given equal consideration as the current generation ($\beta\to 1$).
+That is, in AK models with an infinite planning horizon, giving ultimate consideration to future generations requires the current generation to ultimately reduce their consumption, i.e., $c_{0}^{*}[\infty]\to{\nu}$.
+Alternatively, if we relax the infinite horizon assumption, the initial action for AK models with ultimate consideration for the future generations becomes:
 $c_{0}^{*}[n]=\lim_{\beta\to 1}\frac{1-\beta}{1-\beta^{n}}{A}k_{0}=\frac{Ak_{0}}{n}$
-and we are left with the AK and ZD version of subsistence-proof number of potential generations $n=Ak_{0}/s$.
+and we are left with the AK/ZD version of subsistence-proof size of potential generations, that is, $n=Ak_{0}/\nu$.
 4 Concluding Remarks
-The extent that we take into account of the entire potential generation upon creating the objective function of the population (i.e., the dynasty) is usually determined by the compounding discounting factor.
-Notice that the potential number of intrinsic years that are accounted by compound discounting at a rate of $r$ is as much as $\int_{0}^{\infty}e^{-{r}x}\mathrm{d}x=1/r$.
-As regards to our problem, compound discounting at a rate of $r$ is equivalent to binary discounting at a planning horizon $n=1/r$.
-That is, for example, a generational discount rate of $1$% amounts to 100 generations.
-Binary discounting is flexible in a sense that the planning horizon is variable, so that infinity is not precluded from an option.
-Moreover, it is nondiscriminatory, ensuring that living generations are treated equally.
-With binary discounting, however, an economic growth model requires a finite horizon dynamic optimization approach.
-One important feature of binary discounting is that it hinders future generations that earn below-critical-level utilities.
-If the phase boundary $\log(A\beta)$ for AK setting is negative, the infinite horizon solution trajectory of utility (less the critical level), is forever declining (as shown in the middle-right panel of Figure 3) and eventually surpasses the level of subsistence.
+The extent to which we account for the entire potential generation upon creating the objective function of the population (i.e., the dynasty) is usually determined by the compounding discounting factor.
+The potential number of intrinsic years that are accounted for by compound discounting at a rate of $r$ is as high as $\int_{0}^{\infty}e^{-{r}x}\mathrm{d}x=1/r$.
+With respect to our problem, compound discounting at a rate of $r$ is equivalent to binary discounting (which eliminates consideration for utilities before the planning horizon while fully discounting those that lie beyond it) at a planning horizon $n=1/r$.
+Binary discounting is flexible in the sense that the planning horizon is variable so that infinity is not precluded as an option.
+Moreover, it is nondiscriminatory, ensuring that existing generations are treated in ultimate equity.
+With binary discounting, however, the roundaboutness of production requires a finite horizon dynamic programming approach.
+One important feature of our welfare evaluation on the basis of critical-level utilitarianism is that it hinders future generations that earn below-critical-level utility.
+If the phase boundary $\log(A\beta)$ for the AK setting is negative, the infinite horizon solution trajectory of utility (less the critical level) is forever declining (as shown in the middle-right panel of Figure 3) and eventually surpasses the level of well-being subsistence.
 Binary discounting avoids such repugnancy by limiting the longevity of the dynasty.
-On the other hand, if the phase boundary $\log({A}\Theta)$ for zero (compound) discounting is negative, a finite horizon must be optimal since an infinite horizon objective function approaches a negative infinity.
-Note also that the bottom utilities that lie at the center of the planning horizon (as shown in the middle-right panel of Figure 4), yield increasingly negative contributions to the population value.
-As regards the distribution of consumption among generations, the rate of compound discounting plays the key role.
-That is, the less the rate of discounting is, the more equal the consumptions become.
-Moreover, as the planning horizon expands, the distribution of consumption becomes more uneven in the AK setting cases, whereas vice-versa in a zero discounting setting.
-Our another concern was the optimal initial consumption in light of the well-being of future generations.
-In any case the initial consumption is decreasing in the planning horizon, and ultimately, it will depend on the value of $\beta\theta$.
-The optimal initial consumption for infinite planning horizon will decrease as discount rate approaches zero, and under the AK setting, initial consumption itself approaches zero until it reaches the level of subsistence.
+On the other hand, if the phase boundary $\log B$ for ZD is negative, a finite horizon must be optimal since an infinite horizon objective function approaches negative infinity.
+Note also that the bottom utilities that lie at the center of the planning horizon (as shown in the middle-right panel of Figure 4) yield increasingly negative contributions to the population value.
+With respect to the distribution of consumption among generations, the rate of compound discounting plays a key role.
+That is, the lower the rate of discounting is, the more equal the consumption becomes.
+Moreover, as the planning horizon expands, the distribution of consumption becomes more uneven in an AK setting, whereas vice versa in a ZD setting.
+Another concern was the optimal initial consumption in light of the well-being of future generations.
+In any case, the optimal initial consumption decreases in the planning horizon, and ultimately, it depends on the value of $\beta\theta$.
+If $\beta\theta\to 1$ (that is, with the AK/ZD setting), the optimal initial consumption approaches the level of ultimate sacrifice under an infinite planning horizon, Conversely, any finite consumption of the current generation determines the (finite) planning horizon, i.e., the size of future generations.
 Appendix 1
-Below we put down a production function and its dual unit cost function:
+Below, we write a Cobb–Douglas production function and its dual unit cost function:
 ${Y}=A{K}^{\theta}{L}^{1-\theta},$ $p={B}^{-1}{r}^{\theta}w^{1-\theta}$
-where $p$, $r$, and $w$ denote prices for $Y$, $K$, and $L$, respectively.
-The remaining parameter $A$ is called the productivity and $B$ is called the dual cost function-based productivity.
-Applying Shephard’s lemma to the dual function leads to the followings.
+where $p$, $r$, and $w$ denote prices corresponding to $Y$, $K$, and $L$, respectively.
+The remaining parameter $A$ is referred to as the productivity, and $B$ is the cost function-based productivity.
+Applying Shephard’s lemma on the dual function leads to the following.
 $\frac{\partial p}{\partial r}=\frac{\theta}{B}\left(\frac{r}{w}\right)^{\theta-1}=\frac{K}{Y},$ $\frac{\partial p}{\partial w}=\frac{1-\theta}{B}\left(\frac{r}{w}\right)^{\theta}=\frac{L}{Y}$
-By virtue of these euqations, the marginal product of capital (MPK) can be readily evaluated as follows:
+On the basis of these equations, the marginal product of capital (MPK) can be readily evaluated as follows:
 $\text{MPK}=\frac{\partial Y}{\partial K}=A\theta\left(\frac{K}{L}\right)^{\theta-1}$ $=A\theta^{\theta}(1-\theta)^{1-\theta}\left(\frac{r}{w}\right)^{\theta-1}$
 $A\theta\left(\frac{K}{L}\right)^{\theta-1}$ $=\frac{Y}{K}\theta=B\left(\frac{r}{w}\right)^{\theta-1}$
-where, $r/w\in(0,\infty)$ equals the marginal rate of substitution between capital and labor (MRS). By comparison, we are left with $B=A\theta^{\theta}(1-\theta)^{1-\theta}\equiv A\Theta$. Moreover,
-$A=\left(\frac{\text{MPK}}{\Theta}\right)\left(\frac{r}{w}\right)^{\theta-1}$ $B=\text{MPK}\left(\frac{r}{w}\right)^{\theta-1}$
-On the other hand, recall the breakdown equation of total output (2) and take the partial derivative as follows:
+where $r/w\in(0,\infty)$ is the marginal rate of substitution (MRS) between capital and labor. By comparison, we are left with $B=A\theta^{\theta}(1-\theta)^{1-\theta}$. Moreover,
+$A=\frac{\text{MPK}}{\theta^{\theta}(1-\theta)^{1-\theta}}\left(\frac{r}{w}\right)^{\theta-1},$ $B=\text{MPK}\left(\frac{r}{w}\right)^{\theta-1}$
+On the other hand, we recall the breakdown equation of the total output (2) and take the partial derivative as follows:
 $\frac{\partial Y_{t}}{\partial K_{t}}=\left(\frac{\partial K_{t+1}}{\partial K_{t}}-1\right)+\delta=\rho+\delta=\text{MPK}$ (20)
 where $\rho>0$ denotes the rate of interest. Hence, if $\theta=1$ (AK setting), $A=B=\text{MPK}=\rho+\delta$, and if $\delta=1$ (complete depreciation), it must be the case that $A=B=1+\rho>1$.
-According to the study in section 3.1, we note that condition $A\beta<1$ implies the non-increasing of undiscounted contributions (or utilities less the critical level) with respect to generations.
+From the study in Section 3.1, we note that the condition $A\beta<1$ implies the nonincreasing property of undiscounted contributions (or utilities less the critical level) with respect to generations.
 The knife-edge case where $A\beta=1$ implies that the undiscounted contributions are constant over generations.
-Figure 7 (left) dipicts the set of possible values for $(\theta,A)$ that satisfies $A\beta\leq 1$.
-Also, according to the study in section 3.2, we note that condition $A\Theta<1$ implies the existence of a finite optimal planning horizon.
-Otherwise, if $A\Theta>1$, the optimal planning horizon will be inifinite.
-The knife-edge case where $A\Theta=1$ implies that the optimal planning horizon is indeterminate.
-Figure 7 (right) dipicts the set of possible values for $(\theta,A)$ that satisfies $A\Theta\leq 1$.
-[…]
-Figure 7: The shaded area (mariginal lines not included) corresponds to the set of parameters $(\theta,A)$ where $B=A\theta^{\theta}(1-\theta)^{1-\theta}<1$
+Figure 7 (left) depicts the set of possible values for $(\theta,A)$ that satisfies $A\beta<1$.
+Additionally, from the study in Section 3.2, we note that the condition $B<1$ implies the existence of a finite optimal planning horizon.
+Otherwise, if $B>1$, the optimal planning horizon will be infinite.
+The knife-edge case where $B=1$ implies that the optimal planning horizon is indeterminate.
+Figure 7 (right) depicts the set of possible values for $(\theta,A)$ that satisfies $B<1$.
+Figure 7: The shaded area (marginal lines not included) corresponds to the set of parameters $(\beta,A)$ where $A\beta<1$ (left), and $(\theta,A)$ where $B=A\theta^{\theta}(1-\theta)^{1-\theta}<1$ (right).
 Appendix 2
 Lemma 1.
-The value function of Bellman equation (4) is of the following:
+The value function of the Bellman equation (4) is as follows:
 $\mathcal{V}_{t}\left[k_{t};n\right]$ $=S_{{n}-t}\theta\log(k_{t})+{R}_{t}$
@@
 Proof.
-We show this by induction. First, evaluate the value function at $t={n}$ by (4):
+We show this by induction. First, the value function at $t={n}$ is evaluated via (4):
 $\mathcal{V}_{n}\left[k_{n};{n}\right]=\max_{c_{n}}\left(\log c_{n}+\beta\mathcal{V}_{n+1}[k_{{n}+1}={A}(k_{n})^{\theta}-c_{n};{n}]\right)$
-Since $k_{{n}+1}=0$ by (2.1), $c_{n}={A}(k_{n})^{\theta}$ must be true. Also, $\mathcal{V}_{{n}+1}=0$ must be true for efficiency. Thus, we know that the final maximization is bounded, i.e.,
+Since $k_{{n}+1}=0$ by (2.1), $c_{n}={A}(k_{n})^{\theta}$ must be true. Additionally, $\mathcal{V}_{{n}+1}=0$ must be true for efficiency. Thus, the final maximization is bounded, i.e.,
 $\mathcal{V}_{n}\left[k_{n};{n}\right]=\log c_{n}=\log{A}+\theta\log k_{n}$
-Because $S_{n-n}=S_{0}=1$ and since ${A}$ is a constant, the proposition holds true for $t={n}$.
-Suppose the proposition holds true for $t+1$. Then,
+Because $S_{n-n}=S_{0}=1$ and since ${A}$ is a constant, the lemma holds true for $t={n}$.
+Suppose that the lemma holds true for $t+1$. Then,
 $\mathcal{V}_{t}\left[k_{t};n\right]=$ $\max_{c_{t}}\left(\log c_{t}+\beta\mathcal{V}_{t+1}\left[k_{t+1}={A}(k_{t})^{\theta}-c_{t};n\right]\right)$
 $=$ $\max_{c_{t}}\left(\log c_{t}+\beta\left(S_{{n}-t-1}\theta\log\left({A}(k_{t})^{\theta}-c_{t}\right)+{R}_{t+1}\right)\right)$
-Below are the corresponding first order condition and its solution:
+Below are the corresponding first-order condition and its solution:
 $\frac{1}{c_{t}}-\frac{(\beta\theta)S_{{n}-t-1}}{{A}(k_{t})^{\theta}-c_{t}}=0,$ or, $c_{t}=\frac{{A}(k_{t})^{\theta}}{S_{{n}-t}}$ (21)
-Here, we used $(\beta\theta)S_{{n}-t-1}=\beta\theta+(\beta\theta)^{2}+\cdots+(\beta\theta)^{n-t}=S_{n-t}-1$. By plugging the above solution back into the maximand we arrive at the following result:
+Here, we use $(\beta\theta)S_{{n}-t-1}=\beta\theta+(\beta\theta)^{2}+\cdots+(\beta\theta)^{n-t}=S_{n-t}-1$. By plugging the above solution back into the maximand, we arrive at the following result:
 $\mathcal{V}_{t}[k_{t};n]$ $=\log\left(\frac{{A}(k_{t})^{\theta}}{S_{{n}-t}}\right)+\beta\left(R_{t+1}+S_{n-t-1}\theta\log\left({A}(k_{t})^{\theta}-\frac{{A}(k_{t})^{\theta}}{S_{{n}-t}}\right)\right)$
@@
 $=S_{{n}-t}\theta\log(k_{t})+R_{t}$
-Hence, the lemma. ∎
+Hence, the lemma follows. ∎
 Lemma 2.
-The optimal trajectory of state $k_{t}^{*}[n]$ for Bellman equation (4) is:
-$\ln k_{t}^{*}[n]=\sum_{i=1}^{t}\theta^{t-i}\log\left(\frac{S_{{n}-i}}{S_{{n}-i+1}}{A}\beta\theta\right)+\theta^{t}\log k_{0}$
+The optimal trajectory of state $k_{t}^{*}[n]$ for the Bellman equation (4) is as follows:
+$\log k_{t}^{*}[n]=\sum_{i=1}^{t}\theta^{t-i}\log\left(\frac{S_{{n}-i}}{S_{{n}-i+1}}{A}\beta\theta\right)+\theta^{t}\log k_{0}$
 Proof.
@@
 $k_{t+1}={A}(k_{t})^{\theta}-c_{t}=\frac{S_{{n}-t}-1}{S_{{n}-t}}{A}(k_{t})^{\theta}=\frac{S_{{n}-t-1}}{S_{{n}-t}}{A}\beta\theta(k_{t})^{\theta}$ (22)
-As we apply $t=0$ on the above (22) and by taking the logarithm,
+As we apply $t=0$ to the above (22) and take the logarithm,
 $\log k_{1}=\ln\left(\frac{S_{{n}-1}}{S_{n}}{A}\beta\theta\right)+\theta\log k_{0}$
-we know that the proposition is true for $t=1$.
-Suppose that the proposition is true for $t$. Then, we know by (22) that:
+We know that the lemma is true for $t=1$.
+Suppose that the lemma is true for $t$. We then know by (22) that:
 $\log k_{t+1}^{*}[n]$ $=\log\left(\frac{S_{{n}-t-1}}{S_{{n}-t}}{A}\beta\theta\right)+\theta\log k_{t}^{*}[n]$
@@
 $=\sum_{i=1}^{t+1}\theta^{t+1-i}\log\left(\frac{S_{{n}-i}}{S_{{n}-i+1}}{A}\beta\theta\right)+\theta^{t+1}\log k_{0}$
-That is, the proposition is true for $t+1$. ∎
+which indicates that the lemma is true for $t+1$. Hence, the lemma follows. ∎
 Proposition 1.
-The optimal consumption trajectory $c_{t}^{*}[n]$ for Bellman equation (4) is:
+The optimal consumption trajectory $c_{t}^{*}[n]$ for the Bellman equation (4) is as follows:
 $\log c_{t}^{*}[n]=\log\left(\frac{{A}}{S_{{n}-t}}\right)+\theta\left(\sum_{i=1}^{t}\theta^{t-i}\log\left(\frac{S_{{n}-i}}{S_{{n}-i+1}}{A}\beta\theta\right)+\theta^{t}\log k_{0}\right)$
 Proof.
-Obvious from Proposition 2 and (21). ∎
+This is obvious from Lemma 2 and (21). ∎
 Appendix 3
@@
 $={n(n+1)}/{2}$
-where we used the L’Hôpital’s rule twice. Then, we know that
+where we use L’Hôpital’s rule twice. Then, we know that
 $\lim_{\beta\to 1}\frac{\beta-((1-\beta){n}+1)\beta^{n+1}}{(1-\beta)^{2}}\log(A\beta)=\frac{n(n+1)}{2}\log(A)=0$
+Compliance with Ethical Standards
+The authors declare that they have no conflicts of interest.

### v4

@@
-
+0000-0003-0424-585X
 Abstract
-This paper develops a dynamic optimization framework for evaluating the optimal longevity of a dynasty under a critical-level utilitarian social welfare criterion.
-Unlike standard models that assume an infinite planning horizon and apply discounting to future generations, we endogenize the planning horizon itself and examine the welfare-maximizing number of generations in a roundabout production economy with capital accumulation.
-The model incorporates a well-being threshold that must be met for generational existence to be considered ethically justified, providing an alternative to traditional utilitarian aggregation.
-Using finite-horizon dynamic programming, we derive closed-form consumption trajectories and analyze how technological parameters and intertemporal preferences affect the optimal generational span.
-Our results show that the optimal horizon need not be infinite—even under zero discounting—and that intergenerational equity may require limiting dynasty length.
 […]
-O21 I30 C61
-Keywords:
-Critical-level Utilitarianism Dynamic Programming Optimal Population Intergenerational Equity
 journal:
 1 Introduction
-Economic models of intertemporal planning often assume an infinite horizon, implicitly treating dynasties or economies as perpetually forward-looking entities.
-However, such assumptions can obscure important trade-offs between current and future welfare, particularly when the viability of future generations depends on limited resources and capital dynamics.
-In this paper, we examine a social planner’s problem where the planning horizon—the number of generations considered—is endogenously chosen.
-By adopting a critical-level utilitarian welfare function, we evaluate the contribution of each generation relative to a minimum well-being threshold.
-This framework allows us to study how productivity, capital accumulation, and time preference interact to determine the ethically and economically optimal duration of a dynasty.
-We begin by positing a (classical) unitilitarian social planner whose objective is to indentify a state $x$ that maximizes the sum of the utilities $\mathcal{V}=\sum_{i\in n(x)}\Upsilon_{i}(x)$, which we refer to as the population value, where $n(x)$ denotes the set of individuals and $\Upsilon_{i}(x)$ denotes the lifetime utility (i.e., the level of well-being) of an individual $i$.
-For simplicity, let us assume that all individuals are identical and consume an equal share, denoted by $c$, from a fixed amount of resource $R>0$.
-A utilitarian axiology examines the problem of finding the population size $n=R/c$ that maximizes the population value $\mathcal{V}=\sum_{i=1}^{n}\Upsilon(c_{i})=n\Upsilon(c=R/n)$.
-The utility function $\Upsilon(c)$ is assumed to be monotonically increasing and concave.
-In addition, with respect to the population value $\mathcal{V}=n\Upsilon(c)$, zero utility ($\Upsilon=0$) is equivalent to the absence of life ($n=0$).
-The consumption level $c=\nu$, where $\Upsilon(\nu)=0$, be referred to as the well-being subsistence level, following Dasgupta (2019).
-Owing to the monotonicity of the utility function, $\Upsilon(c)>0$ must hold as long as $c>\nu$ (see Figure 1 (left)).
-It then follows that the domain allowed for $c$ is $(\nu,\infty)$ since $\Upsilon(c)$ represents the level of attainable lifetime utility.
-Figure 1 (left) depicts the optimal solution $c=\omega$, which satisfies the first-order condition for maximizing the population value, i.e., $\Upsilon^{\prime}(\omega)=\Upsilon(\omega)/\omega$.
-Owing to the concavity of the utility function, if the well-being subsistence approaches zero ($\nu\to 0$), as Córdoba (2023) refers to the homothetic case, the optimal consumption also approaches zero ($\omega\to\nu$), while the optimal size of the population approaches infinity ($n^{*}=R/\omega\to\infty$), leading to a conclusion that is considered repugnant (Parfit, 1986).
-Meanwhile, Blackorby et al (2005) introduced a population axiology known as critical-level utilitarianism.
-Within this framework, a social planner aims to determine a set of consumption schedules $c_{i}$ that maximizes the population value $\mathcal{V}=\sum_{i=1}^{n}\left(\Upsilon(c_{i})-\alpha\right)$, where an individual’s contribution to the population value is evaluated relative to the critical utility level $\alpha$ (see Figure 1 (right)).
-That is, an individual $i$’s life contributes positively to the population value if $c_{i}>\kappa=\Upsilon^{-1}(\alpha)$ but negatively otherwise.
-Note that setting $\alpha>0$ prevents the repugnant conclusion even when the underlying utility function is homothetic ($\nu=0$).
-For convenience, we employ a utility function of the form $\Upsilon(c_{i})=\log c_{i}+\alpha$ for analyses so that the value function for the critical-level utilitarian population value in our maximization study becomes $\mathcal{V}=\sum_{i=1}^{n}\left(\Upsilon(c_{i})-\alpha\right)=\sum_{i=1}^{n}\log c_{i}$.
-As we recall our primary idea of incorporating roundabout production into the optimal population framework, let us introduce (discrete) time and assume that the number of individuals per period is fixed to one unit; that is, all (representative) individuals are identified by the period $t$ during which they live.
-[^1] We may therefore refer to $t$ as a generation.
-The population value function then becomes $\mathcal{\mathcal{V}}=\sum_{t=0}^{n}\log c_{t}$.
-Note that this type of value function was widely employed in the previous Ramsey–Cass–Koopmans type of dynastic optimization model (including Nordhaus, 1992; Stern, 2007), except that the future generations were discounted according to the temporal deviation from the present, and typically, the number of potential generations $n$ remains unquestioned (and set to infinity).
-In contrast, in this study, potential generations are accounted for equally, and the number of generations is optimized.
+Standard economic models of optimal growth typically assume an infinite planning horizon.
+This assumption, justified by the mathematical convenience of asymptotic stability or the ethical stance of treating all future generations impartially, implicitly imposes a “perpetual existence” constraint on the dynasty (Ramsey1928).
+While this framework functions well in growing economies, it raises deep ethical concerns when applied to stagnant or declining economies with resource constraints.
+In such settings, the standard discounted utilitarian approach can lead to a “dismal” long-run equilibrium where consumption approaches zero asymptotically (DASGUPTA2019).
+By enforcing indefinite survival, these models may inadvertently justify the perpetuation of lives that are barely worth living, or worse, lives filled with misery—a dynamic analogue to Parfit’s “Repugnant Conclusion” in population ethics (PARFIT).
+This paper revisits the problem of dynastic planning through the lens of Critical-Level Utilitarianism (CLU) (BlackorbyDonaldson1984; BBD2005).
+Unlike standard utilitarianism, which sums utilities regardless of their sign, CLU introduces a normative threshold of well-being (the critical level).
+An individual’s existence contributes to social welfare only if their utility exceeds this critical level; otherwise, it is considered a negative contribution.
+By applying this ethical principle to a dynamic setting, we pose a question that is often excluded from standard growth theory: Under what economic conditions is it ethically optimal for a dynasty to continue, and when is its planned termination the morally justified conclusion of history?
+Methodologically, our approach establishes a structural isomorphism between static population ethics and dynamic growth theory.
+In the tradition of optimal population theory initiated by Samuelson1975, a social planner typically chooses the optimal population size $N$ to maximize social welfare under resource constraints.
+In our dynamic framework, we reinterpret this problem by viewing the number of generations $N$ (longevity) as the choice variable.
+This mapping reveals that the “Repugnant Conclusion” (a large population with low quality of life) and “Infinite Misery” (an infinite dynasty with low quality of life) are two sides of the same ethical coin.
+Consequently, our finite-horizon solution implies a conscious decision to cease reproduction, rather than a failure of time consistency.
+It represents an altruistic decision to prevent the creation of future generations whose lives would fall below the critical level of well-being.
+We acknowledge that other axiomatic frameworks, such as Rank-Discounted Utilitarianism (RDU), have been proposed to avoid the Repugnant Conclusion (AsheimZuber2014).
+RDU prevents the conclusion by strictly discounting the value of additional lives.
+However, this approach can lead to the opposite extreme—an “anti-population bias” or “reverse repugnant conclusion”—where a small population with high utility is preferred over a larger population that is also well-off.
+Furthermore, in a dynamic context, RDU implies that the welfare weight of future generations depends on their relative rank in history.
+This dependency destroys the recursive structure of the optimization problem and leads to time inconsistency.
+In contrast, CLU provides a transparent, absolute threshold for “lives worth living” that preserves both intergenerational equity (for those above the threshold) and the tractability of dynamic programming.
 The remainder of the paper proceeds as follows.
-In the following section, we first describe the model of roundabout production on the basis of Cobb–Douglas technology.
-We then define our main problem in the form of finite horizon dynamic programming, where the objective function comprises a series of Bernoullian contributions, and analytically derive a general solution.
-In Section 3, we introduce two broad settings of parameters to simplify the general solution and allow us to study the optimal trajectory of the variables in a numerical fashion.
-The first setting is the AK setting, where the output elasticity of capital is set to unity.
-The second is the zero discounting (ZD) setting, where the discount factor is set to unity.
-We also study the peculiarity of the two models with respect to intergenerational equity and the optimal initial action of consumption.
-Section 4 concludes the paper.
+Section 2 describes the model of roundabout production using Cobb–Douglas technology and defines the main problem as finite horizon dynamic programming.
+We analytically derive the optimal consumption path and the optimal longevity $N^{*}$.
+In Section 3, we analyze the properties of the solution under two parameter settings: the AK setting (unity output elasticity of capital) and the Zero Discounting (ZD) setting (intergenerational equity).
+We show that under specific conditions—particularly when productivity is low—the ethically optimal policy is to choose a finite horizon, suggesting that sustainability does not always imply indefinite continuation.
+Section 4 concludes with discussions on the ethical implications of our findings for modern demographic trends.
+2 Model
+2.1 Critical-Level Preferences
+Before describing the production technology, we first specify the social planner’s objective function.
+We posit a social planner whose objective is to maximize the population value based on Critical-Level Utilitarianism (CLU) (BlackorbyDonaldson1984; BBD2005).
+Let $\Upsilon(c)$ denote the lifetime utility (well-being) of an individual consuming a constant level of resources $c$.
+The planner aims to determine the consumption path and the planning horizon (number of generations) to maximize the following social welfare function:
+$\mathcal{V}=\sum_{t=0}^{N}\left(\Upsilon(c_{t})-\alpha\right)$ (1)
+where $\alpha$ represents the critical level of utility.
+An individual’s life contributes positively to social welfare if and only if their utility exceeds this critical threshold $\alpha$; otherwise, it reduces social welfare.
+For analytical tractability and to capture the essential trade-offs, we employ a logarithmic utility function of the form $\Upsilon(c)=\log c+\alpha$.
+This functional form is widely employed in optimal growth literature, including climate change economics (Nordhaus1992; Stern2007), which facilitates comparisons with standard benchmarks.
+Under this specification, the contribution of generation $t$ to the population value becomes:
+$\Upsilon(c_{t})-\alpha=(\log c_{t}+\alpha)-\alpha=\log c_{t}$ (2)
+Consequently, the objective function simplifies to $\mathcal{V}=\sum_{t=0}^{N}\log c_{t}$.
+We define the well-being subsistence level $\nu$ as the consumption level where utility is exactly zero, i.e., $\Upsilon(\nu)=0$.
+In our logarithmic specification, this corresponds to $\nu=e^{-\alpha}$.
+As illustrated in Figure 1, the domain of permissible consumption is $c>\nu$.
+The critical level $\alpha$ thus serves as a “buffer” above the biological subsistence level.
+The planner will only choose to extend the dynasty if the consumption level allows for a utility strictly greater than the critical level (i.e., $\log c_{t}>0$, or $c_{t}>1$ in normalized terms), ensuring that each generation enjoys a life worth living.
 Figure 1: Left: Utility function of an individual.
@@
 In all cases, the domain of permissible consumption levels $c$ for existence is $c>\nu$.
-2 Model
 2.2 Roundabout production
@@
 where $Y_{t}$ denotes the economy’s output, $K_{t}$ denotes capital, and $L_{t}$ denotes labor, all of which are effective during period $t$.
-For the relevant parameters, $10<\theta<$ 1 denotes the output elasticity of capital.
+For the relevant parameters, $0<\theta<1$ denotes the output elasticity of capital.
 The level of technology is indicated by productivity, denoted by $A$, which is assumed to be fixed throughout the considered time span.
@@
 Additionally, $y_{t}=Y_{t}/L$ and $c_{t}=C_{t}/L$ denote per capita output and consumption, respectively, in period $t$.
+The social planner aims to maximize the critical-level utilitarian population value $\mathcal{\mathcal{V}}$ as previously specified, subject to the state transition function (5), i.e.,
 […]
+[…]
+given the initial state $k_{0}$, we assume complete depreciation $\delta=1$ and introduce the discount factor $\beta\leq 1$.
+We let $\beta\to 1$ for a (critical-level) utilitarian assessment.
+[^2] The discount factor $\beta$ is otherwise referred to as the rate of time preference.
+In the context of dynastic optimization, $\beta$ may be referred to as the rate of generational preference (of the population).
+The optimal consumption path $(c_{0}^{*},c_{1}^{*},\cdots,c_{N}^{*})$ is clearly dependent on the planning horizon $N$.
+We therefore solve the above problem hierarchically.
+That is, we first solve for the optimal consumption path given $N$ via finite horizon dynamic programming and obtain the population value function with respect to the planning horizon $\mathcal{\mathcal{V}}[N]$; therefore, we search for the optimal planning horizon $N^{*}$.
 2.3 Finite horizon dynamic programming
 The primary problem here is to solve for an optimal consumption path given a planning horizon.
-The Bellman equation of the problem is as follows:444 Note that the objective function of the primary problem (2.1) is given at $t=0$, i.e., $\mathcal{V}[n]=\mathcal{V}_{0}[k_{0};n]$.
+The Bellman equation of the problem is as follows: [^3] Note that the objective function of the primary problem (6a) is given at $t=0$, i.e., $\mathcal{V}[N]=\mathcal{V}_{0}[k_{0};N]$.
 $\mathcal{\mathcal{V}}_{t}\left[k_{t};\ N\right]=\max_{c_{t}}\left(\log c_{t}+\beta\mathcal{\mathcal{V}}_{t+1}\left[k_{t+1}={A}(k_{t})^{\theta}-c_{t};\ N\right]\right)$ (7)
@@
 Table 1: Parameters applied in various cases.
-case ${A}$ $\beta$ $\theta$ $\log({A}\beta)$ $\log B$ $n^{*}$ $\mathcal{V}[n^{*}]$ $\mathcal{V}[\infty]$
-I $1.012$ $0.992$ $1$ $+$ $\infty$ $84.7$
-II $1.01$ $0.992$ $1$ $+$ $95$ $60.0$
-III $1/\beta$ $0.992$ $1$ 00 $73$ $55.6$
-IV $1.005$ $0.992$ $1$- $58$ $51.0$
-V $1.05$ $1$ $0.992$ $+$ $\infty$ +$\infty$+ ∞
-VI $1.2$ $1$1 00 $(281)$ $41.7$
-VII $1.05$ $1$ $0.991$- $117$$\infty$- ∞
-VIII $1$ $1$ $1$ 00 00 $54$ $55.2$$\infty$- ∞
-IX $1/\beta$ $0.992$2 00 00 $53$ $49.1$ $-1843.2$
+case A{A} $\beta$ $\theta$ $\log({A}\beta)$ $\log B$ $N^{*}$ $\mathcal{V}[N^{*}]$ $\mathcal{V}[\infty]$
+I $1.012$ $0.992$ 11 ++ $\infty$ $84.7$
+II $1.01$ $0.992$ 11 ++ $95$ $60.0$
+III $1/\beta$ $0.992$ 11 0 $73$ $55.6$
+IV $1.005$ $0.992$ 11 - $58$ $51.0$
+V $1.05$ 11 $0.992$ ++ $\infty$ $\infty$
+VI $1.2$ 11 *1 0 $(281)$ $41.7$
+VII $1.05$ 11 $0.991$ - $117$ -$\infty$
+VIII 11 11 11 0 0 $54$ $55.2$ -$\infty$
+IX $1/\beta$ $0.992$ *2 0 0 $53$ $49.1$ 1843.2-1843.2
   • Note:
@@
   • *1
-    $\theta\approx 0.955392$ where 1.2($1{1.2}(1-\theta)^{1-\theta}\theta^{\theta}=B=$ 1.
+    $\theta\approx 0.955392$ where ${1.2}(1-\theta)^{1-\theta}\theta^{\theta}=B=1$.
   • *2
@@
 3.1 AK setting
-The AK model, which was formally developed by Frankel (1962), is one of the simplest fundamental models of endogenous growth.
+The AK model, which was formally developed by Frankel, is one of the simplest fundamental models of endogenous growth.
 Here, we employ this production model to study the population value function that discounts future generations to determine whether $N\to\infty$ is an optimal policy.
@@
 $=\frac{\beta-((1-\beta){N}+1)\beta^{{N}+1}}{(1-\beta)^{2}}\log({A}\beta)+\frac{1-\beta^{{N}+1}}{1-\beta}\log\left(\frac{1-\beta}{1-\beta^{{N}+1}}{A}k_{0}\right)$ (13)
-For the sake of the analysis, let us take the derivative with respect to ${n}$.
+For the sake of the analysis, let us take the derivative with respect to N{N}.
 $\frac{\mathrm{d}\mathcal{V}[N]}{\mathrm{d}N}=\lambda\left(\gamma-{N}\log({A}\beta)+\log\left(1-\beta^{N+1}\right)\right)$ (14)
@@
 In AK models with complete capital depreciation, $\theta=\delta=1$ leads to ${A}=1+\rho$.
-Therefore, assuming that ${A}>1$ is relevant in this setting.555 Note also that $A\beta=(\delta+\rho)\beta$ indicates the discrepancy between the gross interest rate and generational preference rate of the population.
+Therefore, assuming that A>1{A}>1 is relevant in this setting.
+[^4] Note also that $A\beta=(\delta+\rho)\beta$ indicates the discrepancy between the gross interest rate and generational preference rate of the population.
+Figure 2: Left: The solid lines represent $g[N]$, as defined in (15), for cases I–IV with clear correspondences, i.e., the steepest slope corresponds to case I, while case IV corresponds to the scenario with a negative slope.
+The dashed line represents $f[N]$, which is also defined in (15).
+Right: The plots show population value functions $\mathcal{V}[N]$ for cases I–IV with obvious correspondences.
 Figure 2 (left) depicts the functions $f[N]$ and $g[N]$ under different parameters for cases I–IV.
 Clearly, $f[N]$ is a linear function whose slope is $\log({A}\beta)$, and $g[N]$ monotonously increases and approaches zero, i.e., $g[\infty]=0$.
-Here, we fix the discount factor $\beta$ and differentiate the productivity ${A}$ at four different levels.
+Here, we fix the discount factor $\beta$ and differentiate the productivity A{A} at four different levels.
 As long as ${A}\leq 1/\beta$, so that $\log({A}\beta)\leq 0$, i.e., the slope is zero or negative, $f[N]$ will intersect with $g[N]$ at a single point, and the population value function $\mathcal{V}[N]$ has a single peak.
@@
 Figure 2 (right) depicts the population value functions for cases I–IV.
-Figure 2: Left: The solid lines represent $g[n]$, as defined in (12), for cases I–IV with clear correspondences, i.e., the steepest slope corresponds to case I, while case IV corresponds to the scenario with a negative slope.
-The dashed line represents $f[n]$, which is also defined in (12).
-Right: The plots show population value functions $\mathcal{V}[n]$ for cases I–IV with obvious correspondences.
 To visualize the optimal trajectory of the variables in various situations, we specify them here in the form of functionals.
@@
 Similarly, the bottom row of Figure 3 displays the optimal trajectories of capital intensity, based on (18), for planning horizons $N=200,400,600$ for cases I, III, and IV (from left to right).
-Figure 3: Population value function (top), trajectories of undiscounted contribution (middle) and capital intensity (bottom), for AK setting cases I (left), III (center), and IV (right).
-The parameters corresponding to each of the cases are given in Table 1.
+Figure 3: Population value function (top), seleted trajectories of undiscounted contribution (middle) and capital intensity (bottom), for AK setting cases I (left), III (center), and IV (right).
+The trajectories are selected for $N=200,400,600$.
 3.2 ZD setting
@@
 $=\frac{\log\left({A}(1-\beta\theta)^{1-\theta}(\beta\theta)^{\theta}\right)}{1-\theta}+\theta^{t}\left(\log(k_{0})^{\theta}-\log({A}\beta\theta)^{\frac{\theta}{1-\theta}}\right)$ (20)
-where the optimal trajectory for undiscounted contributions is geometrically convergent.
-We can then apply $\beta=1$ to arrive at the following zero discount verification of the optimal trajectory for undiscounted contributions:
+where the optimal trajectory for undiscounted contributions is geometrically convergent. We can then apply $\beta=1$ to arrive at the following:
 $\log c_{t}^{*}[\infty]=\frac{\log\left({A}(1-\theta)^{1-\theta}\theta^{\theta}\right)}{1-\theta}+\theta^{t}\left(\log(k_{0})^{\theta}-\log({A}\theta)^{\frac{\theta}{1-\theta}}\right)$
@@
 The case where $B=1$ is the knife-edge case in which the planning horizon $N$ does not matter (beyond a certain length) in maximizing the population value.
-Case V corresponds to $B>1$, and $n^{*}=\infty$; case VII corresponds to $B<1$, and $n^{*}\ll\infty$; and case VI corresponds to $B=1$, and $n^{*}$ is any number greater than $281$.777 The number is subject to decimal rounding.
+Case V corresponds to $B>1$, and $N^{*}=\infty$; case VII corresponds to $B<1$, and $N^{*}\ll\infty$; and case VI corresponds to $B=1$, and $N^{*}$ is any number greater than 281281.666The number is subject to decimal rounding.
 Figure 4 depicts the population value function $\mathcal{V}[N]$, along with the optimum trajectory of undiscounted contribution $\log c_{t}^{*}[N]$ and the capital intensity $k_{t}^{*}[N]$, given a planning horizon $N$, for sample ZD settings V, VI, and VII.
@@
 By the first-order condition $\frac{\partial\mathcal{V}[N]}{\partial N}=0$, the optimal planning horizon is evaluated as $N^{*}=\exp(-1+\log k_{0})-1\approx 54.182$.
-Figure 4: Population value function (top), trajectories of undiscounted contribution (middle) and capital intensity (bottom), for ZD setting cases V (left), VI (center), and VII (right).
+Figure 4: Population value function (top), selected trajectories of undiscounted contribution (middle) and capital intensity (bottom), for ZD setting cases V (left), VI (center), and VII (right).
+The trajectories are selected for $N=200,400,600$.
 Note that the population value function of the top-left panel is increasing indefinitely.
@@
 4 Concluding Remarks
-The extent to which we account for the entire potential generation upon creating the objective function of the population (i.e., the dynasty) is usually determined by the compounding discounting factor.
-The potential number of intrinsic years that are accounted for by compound discounting at a rate of $r$ is as high as $\int_{0}^{\infty}e^{-{r}x}\mathrm{d}x=1/r$.
-With respect to our problem, compound discounting at a rate of $r$ is equivalent to binary discounting (which eliminates consideration for utilities before the planning horizon while fully discounting those that lie beyond it) at a planning horizon $n=1/r$.
-Binary discounting is flexible in the sense that the planning horizon is variable so that infinity is not precluded as an option.
-Moreover, it is nondiscriminatory, ensuring that existing generations are treated in ultimate equity.
-With binary discounting, however, the roundaboutness of production requires a finite horizon dynamic programming approach.
-One important feature of our welfare evaluation on the basis of critical-level utilitarianism is that it hinders future generations that earn below-critical-level utility.
-If the phase boundary $\log(A\beta)$ for the AK setting is negative, the infinite horizon solution trajectory of utility (less the critical level) is forever declining (as shown in the middle-right panel of Figure 3) and eventually surpasses the level of well-being subsistence.
-Binary discounting avoids such repugnancy by limiting the longevity of the dynasty.
-On the other hand, if the phase boundary $\log B$ for ZD is negative, a finite horizon must be optimal since an infinite horizon objective function approaches negative infinity.
-Note also that the bottom utilities that lie at the center of the planning horizon (as shown in the middle-right panel of Figure 4) yield increasingly negative contributions to the population value.
-With respect to the distribution of consumption among generations, the rate of compound discounting plays a key role.
-That is, the lower the rate of discounting is, the more equal the consumption becomes.
-Moreover, as the planning horizon expands, the distribution of consumption becomes more uneven in an AK setting, whereas vice versa in a ZD setting.
-Another concern was the optimal initial consumption in light of the well-being of future generations.
-In any case, the optimal initial consumption decreases in the planning horizon, and ultimately, it depends on the value of $\beta\theta$.
-If $\beta\theta\to 1$ (that is, with the AK/ZD setting), the optimal initial consumption approaches the level of ultimate sacrifice under an infinite planning horizon, Conversely, any finite consumption of the current generation determines the (finite) planning horizon, i.e., the size of future generations.
+This paper has investigated the optimal longevity of a dynasty under a Critical-Level Utilitarian framework.
+By treating the planning horizon as an endogenous choice variable within a dynamic programming setting, we derived closed-form analytical solutions that link economic fundamentals to the ethical duration of a society.
+Our analysis highlights two distinct regimes determined by productivity and time preference.
+In the AK setting with discounting, the sign of $\log(A\beta)$ serves as the conservative boundary determining whether the optimal horizon is finite or infinite.
+Similarly, in the Zero Discounting (ZD) setting, which prioritizes intergenerational equity, the sign of the cost-based productivity $\log(B)$ dictates the optimal longevity.
+These results demonstrate that under low productivity or strict ethical constraints, a finite horizon is not an anomaly but a structurally optimal response to avoid negative social welfare.
+Our findings also reveal a striking contrast in intergenerational inequality between the two settings.
+In the AK setting, a longer planning horizon is associated with greater inequality across generations, although a higher discount factor $\beta$ mitigates this effect.
+Conversely, in the ZD setting, extending the horizon tends to reduce inequality, as the burden of capital accumulation is shared more evenly.
+However, a higher output elasticity of capital $\theta$ in the ZD model exacerbates long-term inequality.
+These observations suggest that the pursuit of longevity imposes different distributional costs depending on the underlying technology and ethical discount rates.
+A key theoretical insight emerges from the behavior of the initial consumption $c_{0}$ as we approach the limit of perfect intergenerational impartiality ($\beta\to 1$).
+In both models, a higher $\beta$ generally requires the current generation to reduce consumption to support a longer dynasty.
+In the limiting case of the infinite-horizon AK model with $\beta\to 1$ (and $\theta=1$), the optimal initial consumption approaches zero (or the subsistence level).
+This implies that if the future holds infinite potential, the current generation is ethically compelled to make the ultimate sacrifice for the sake of an eternal dynasty.
+In contrast, in a stagnant economy where such sacrifice yields diminishing returns, our model prescribes termination.
+Thus, our framework illuminates an ethical asymmetry: it demands immense endurance when the future is promising, but offers the “mercy” of termination when the future promises only misery.
+Methodologically, our approach ensures time consistency through the recursive structure of dynamic programming.
+Unlike ad hoc termination rules, the finite horizon derived here is the result of a consistent optimization process where the continuation value falls below the critical threshold.
+We acknowledge, however, that our model relies on a logarithmic utility function and a fixed population size per generation to maintain analytical tractability.
+Introducing more general preferences (such as CRRA utility), endogenous population size, or environmental constraints would likely preclude closed-form solutions and require numerical approaches.
+Exploring these complex dynamics remains a promising avenue for future research.
+Finally, our results offer a normative perspective on the demographic trends observed in modern mature economies.
+The choice to limit the number of offspring—or in our macroeconomic interpretation, to limit the longevity of the dynasty—is often viewed as a failure of vitality.
+However, from the perspective of Critical-Level Utilitarianism, this may be interpreted as an altruistic termination.
+If the current generation anticipates that economic or environmental constraints will condemn future generations to lives barely worth living, choosing a finite horizon is a rational and ethical decision to prevent the creation of suffering.
 Appendix 1
@@
 $\frac{\partial Y_{t}}{\partial K_{t}}=\left(\frac{\partial K_{t+1}}{\partial K_{t}}-1\right)+\delta=\rho+\delta=\text{MPK}$ (23)
-where $\rho>0$ denotes the rate of interest. Hence, if $\theta=1$ (AK setting), $A=B=\text{MPK}=\rho+\delta$, and if $\delta=1$ (complete depreciation), it must be the case that $A=B=1+\rho>1$.
+where $\rho>0$ denotes the rate of interest.
+Hence, if $\theta=1$ (AK setting), $A=B=\text{MPK}=\rho+\delta$, and if $\delta=1$ (complete depreciation), it must be the case that $A=B=1+\rho>1$.
 From the study in Section 3.1, we note that the condition $A\beta<1$ implies the nonincreasing property of undiscounted contributions (or utilities less the critical level) with respect to generations.
@@
 Proof.
-We show this by induction. First, the value function at $t={n}$ is evaluated via (4):
+We show this by induction. First, the value function at t=Nt={N} is evaluated via (7):
 $\mathcal{V}_{N}\left[k_{N};{N}\right]=\max_{c_{N}}\left(\log c_{N}+\beta\mathcal{V}_{N+1}[k_{{N}+1}={A}(k_{N})^{\theta}-c_{N};{N}]\right)$
-Since $k_{{n}+1}=0$ by (2.1), $c_{n}={A}(k_{n})^{\theta}$ must be true. Additionally, $\mathcal{V}_{{n}+1}=0$ must be true for efficiency. Thus, the final maximization is bounded, i.e.,
+Since $k_{{N}+1}=0$ by (6b), $c_{N}={A}(k_{N})^{\theta}$ must be true. Additionally, $\mathcal{V}_{{N}+1}=0$ must be true for efficiency. Thus, the final maximization is bounded, i.e.,
 $\mathcal{V}_{N}\left[k_{N};{N}\right]=\log c_{N}=\log{A}+\theta\log k_{N}$
-Because $S_{n-n}=S_{0}=1$ and since ${A}$ is a constant, the lemma holds true for $t={n}$.
-Suppose that the lemma holds true for $t+1$. Then,
+Because $S_{N-N}=S_{0}=1$ and since A{A} is a constant, the lemma holds true for t=Nt={N}. Suppose that the lemma holds true for $t+1$. Then,
 $\mathcal{V}_{t}\left[k_{t};N\right]=$ $\max_{c_{t}}\left(\log c_{t}+\beta\mathcal{V}_{t+1}\left[k_{t+1}={A}(k_{t})^{\theta}-c_{t};N\right]\right)$
@@
 Proof.
-We show this by induction. By plugging (21) into (2.1), we obtain:
+We show this by induction. By plugging (24) into (6b), we obtain:
 $k_{t+1}={A}(k_{t})^{\theta}-c_{t}=\frac{S_{{N}-t}-1}{S_{{N}-t}}{A}(k_{t})^{\theta}=\frac{S_{{N}-t-1}}{S_{{N}-t}}{A}\beta\theta(k_{t})^{\theta}$ (25)
@@
 $\log k_{1}=\ln\left(\frac{S_{{N}-1}}{S_{N}}{A}\beta\theta\right)+\theta\log k_{0}$
-We know that the lemma is true for $t=1$.
-Suppose that the lemma is true for $t$. We then know by (22) that:
+We know that the lemma is true for $t=1$. Suppose that the lemma is true for $t$. We then know by (25) that:
 $\log k_{t+1}^{*}[N]$ $=\log\left(\frac{S_{{N}-t-1}}{S_{{N}-t}}{A}\beta\theta\right)+\theta\log k_{t}^{*}[N]$

### v5

@@
 Standard economic models of optimal growth typically assume an infinite planning horizon.
-This assumption, justified by the mathematical convenience of asymptotic stability or the ethical stance of treating all future generations impartially, implicitly imposes a “perpetual existence” constraint on the dynasty (Ramsey1928).
+This assumption, justified by the mathematical convenience of asymptotic stability or the ethical stance of treating all future generations impartially, implicitly imposes a “perpetual existence” constraint on the dynasty (Ramsey, 1928).
 While this framework functions well in growing economies, it raises deep ethical concerns when applied to stagnant or declining economies with resource constraints.
-In such settings, the standard discounted utilitarian approach can lead to a “dismal” long-run equilibrium where consumption approaches zero asymptotically (DASGUPTA2019).
-By enforcing indefinite survival, these models may inadvertently justify the perpetuation of lives that are barely worth living, or worse, lives filled with misery—a dynamic analogue to Parfit’s “Repugnant Conclusion” in population ethics (PARFIT).
-This paper revisits the problem of dynastic planning through the lens of Critical-Level Utilitarianism (CLU) (BlackorbyDonaldson1984; BBD2005).
+In such settings, the standard discounted utilitarian approach can lead to a “dismal” long-run equilibrium where consumption approaches zero asymptotically (Dasgupta, 2019).
+By enforcing indefinite survival, these models may inadvertently justify the perpetuation of lives that are barely worth living, or worse, lives filled with misery—a dynamic analogue to Parfit’s “Repugnant Conclusion” in population ethics (Parfit, 1986).
+This paper revisits the problem of dynastic planning through the lens of Critical-Level Utilitarianism (CLU) (Blackorby and Donaldson, 1984; Blackorby et al., 2005).
 Unlike standard utilitarianism, which sums utilities regardless of their sign, CLU introduces a normative threshold of well-being (the critical level).
@@
 By applying this ethical principle to a dynamic setting, we pose a question that is often excluded from standard growth theory: Under what economic conditions is it ethically optimal for a dynasty to continue, and when is its planned termination the morally justified conclusion of history?
-Methodologically, our approach establishes a structural isomorphism between static population ethics and dynamic growth theory.
-In the tradition of optimal population theory initiated by Samuelson1975, a social planner typically chooses the optimal population size $N$ to maximize social welfare under resource constraints.
-In our dynamic framework, we reinterpret this problem by viewing the number of generations $N$ (longevity) as the choice variable.
+Methodologically, our approach establishes a structural isomorphism between the spatial and temporal dimensions of optimal population theory.
+In the tradition initiated by Samuelson (1975), dynamic models of optimal population typically assume an infinite planning horizon and focus on choosing the optimal population growth rate (i.e., the spatial size of cohorts) to maximize social welfare.
+Recently, Renström and Spataro (2011, 2019) extended this framework to (Relative) CLU.
+Yet, across this literature, the assumption of perpetual existence remains unquestioned.
+Our paper complements this literature by inverting the paradigm: rather than optimizing the horizontal size of each generation over an infinite horizon, we fix the cohort size and treat the vertical length of the dynasty—the number of generations $N$ (longevity)—as the endogenous choice variable.
 This mapping reveals that the “Repugnant Conclusion” (a large population with low quality of life) and “Infinite Misery” (an infinite dynasty with low quality of life) are two sides of the same ethical coin.
@@
 It represents an altruistic decision to prevent the creation of future generations whose lives would fall below the critical level of well-being.
-We acknowledge that other axiomatic frameworks, such as Rank-Discounted Utilitarianism (RDU), have been proposed to avoid the Repugnant Conclusion (AsheimZuber2014).
+We acknowledge that other axiomatic frameworks, such as Rank-Discounted Utilitarianism (RDU), have been proposed to avoid the Repugnant Conclusion (Asheim and Zuber, 2014).
 RDU prevents the conclusion by strictly discounting the value of additional lives.
@@
 Before describing the production technology, we first specify the social planner’s objective function.
-We posit a social planner whose objective is to maximize the population value based on Critical-Level Utilitarianism (CLU) (BlackorbyDonaldson1984; BBD2005).
+We posit a social planner whose objective is to maximize the population value based on Critical-Level Utilitarianism (CLU) (Blackorby and Donaldson, 1984; Blackorby et al., 2005).
 Let $\Upsilon(c)$ denote the lifetime utility (well-being) of an individual consuming a constant level of resources $c$.
@@
 For analytical tractability and to capture the essential trade-offs, we employ a logarithmic utility function of the form $\Upsilon(c)=\log c+\alpha$.
-This functional form is widely employed in optimal growth literature, including climate change economics (Nordhaus1992; Stern2007), which facilitates comparisons with standard benchmarks.
+This functional form is widely employed in optimal growth literature, including climate change economics (Nordhaus, 1992; Stern, 2007), which facilitates comparisons with standard benchmarks.
 Under this specification, the contribution of generation $t$ to the population value becomes:
@@
 3.1 AK setting
-The AK model, which was formally developed by Frankel, is one of the simplest fundamental models of endogenous growth.
+The AK model, which was formally developed by Frankel (1962), is one of the simplest fundamental models of endogenous growth.
 Here, we employ this production model to study the population value function that discounts future generations to determine whether $N\to\infty$ is an optimal policy.
@@
 $=\frac{\beta-((1-\beta){N}+1)\beta^{{N}+1}}{(1-\beta)^{2}}\log({A}\beta)+\frac{1-\beta^{{N}+1}}{1-\beta}\log\left(\frac{1-\beta}{1-\beta^{{N}+1}}{A}k_{0}\right)$ (13)
-For the sake of the analysis, let us take the derivative with respect to N{N}.
+For the sake of the analysis, let us treat the planning horizon $N$ as a continuous variable and take the derivative with respect to $N$.
 $\frac{\mathrm{d}\mathcal{V}[N]}{\mathrm{d}N}=\lambda\left(\gamma-{N}\log({A}\beta)+\log\left(1-\beta^{N+1}\right)\right)$ (14)
@@
 Methodologically, our approach ensures time consistency through the recursive structure of dynamic programming.
+[^7] Formally, suppose that at some intermediate period $t<N^{*}$, the planner re-optimizes and chooses a different horizon $N^{\#}\neq N^{*}$ (and a corresponding consumption path) that yields a higher continuation value than the original plan.
+This would imply that the original total value could have been improved by switching to this new path at $t$, contradicting the optimality of the original plan $(c_{0}^{*},\dots,c_{N^{*}}^{*})$.
+Therefore, the optimal horizon $N^{*}$ determined at $t=0$ remains optimal for any subsequent generation $t$.
 Unlike ad hoc termination rules, the finite horizon derived here is the result of a consistent optimization process where the continuation value falls below the critical threshold.